### v1

# Closed mean curvature flows with asymptotically conical singularities

Abstract.
In this paper, we prove that for any asymptotically conical self-shrinker, there exists a closed hypersurface such that the mean curvature flow starting from it develops a singularity modeled on the given shrinker. The main technique is the Ważewski box argument, used by Stolarski [Sto] in the proof of the corresponding theorem in the Ricci flow case.
1. Introduction
We consider a family of immersions $\mathbf{x}\colon M^{n}\times I\to\mathbb{R}^{n+k}$ such that $\partial_{t}\mathbf{x}=\mathbf{H}$ where $\mathbf{H}$ is the mean curvature vector of the immersed submanifold $M_{t}=\mathbf{x}(M,t).$ This family of immersions is called a mean curvature flow. Understanding the singularity formations in a mean curvature flow is a fundamental question. More precisely, given a mean curvature flow $M_{t}$ developing a spacetime singularity at $(\textbf{0},0)\in\mathbb{R}^{n+k}\times\mathbb{R},$ we can consider a sequence of rescaled mean curvature flows $M^{i}_{t}:=\lambda_{i}M_{\lambda_{i}^{-2}t}$ using any sequence $\lambda_{i}\to\infty.$ Based on Huisken’s monotonicity [H90] and Brakke’s compactness [B78], White [W94] and Ilmanen [I95] proved that $M_{t}^{i}$ weakly converges to a self-similar limit flow, called a tangent flow, after replacing $\lambda_{i}$’s with a subsequence. When the limiting flow is smooth, it is generated by a self-shrinker, a submanifold satisfying $\textbf{H}=-\frac{x^{\perp}}{2}.$ Uniqueness of tangent flows, a fundamental issue for singularity analysis, has been discussed for certain shrinkers [S14, CM15, CM19, Z20, CS, LSS, LZ, CSSZ]. For related references on uniqueness of tangent cones of minimal submanifolds, one can see [AA, S83, H97, S08, CM14].
In this paper, we focus on singularities modeled on asymptotically conical self-shrinkers of codimension one. An asymptotically conical self-shrinker is a self-shrinker that is asymptotic to a regular cone at infinity. They are especially important in understanding two-dimensional self-shrinkers; cf. [W16]. Moreover, given any cone in $\mathbb{R}^{n+k}$, there exists at most one asymptotically conical self-shrinker asymptotic to it even in the case that the codimension is higher than one [W14, Kh]. Also, asymptotically conical self-shrinkers as tangent flows are unique, proven by Chodosh-Schulze [CS] in the hypersurface case and the authors [LZ] in the higher codimensional case.
A fundamental question on the singularity formation of mean curvature flows is whether there is a closed hypersurface that develops a singularity modeled on an asymptotically conical shrinker. Note that an asymptotically conical shrinker is always non-compact; so is the self-shrinking solution modeled on it. In general, it is not easy to determine the precise type of singularities developed along a mean curvature flow. Thus, although there are some examples of asymptotically conical shrinkers [N14, KKM], a construction of a closed hypersurface flow developing such a singularity has not been given before, though it is expected. In this paper, we prove the following theorem and confirm this expectation.
[…]
Given an asymptotically conical shrinker $\Sigma^{n}\subset\mathbb{R}^{n+1}$, there exists a closed embedded hypersurface $M\subset\mathbb{R}^{n+1}$ such that the mean curvature flow $M_{t}$ starting from $M$ develops a type I singularity at time $1$ at the origin modeled on $\Sigma$.
This confirms the belief in the existence of closed mean curvature flows having asymptotically conical tangent flows. It then has some applications based on earlier works. For example, Sun-Xue [SX2] proved that given a closed mean curvature flow with an asymptotically conical singularity, after a generic initial perturbation, the perturbed mean curvature flow avoids the asymptotically conical singularity. Their theorem suggests that it is hard to see an asymptotically conical singularity in closed mean curvature flows. In fact, it is expected and has been proven in the surface case (cf. [CCMS, CCS, SX1, SX2, BK23]) that all non-cylindrical singularities are not generic.
The proof of Theorem 1 follows the outline of the argument of the corresponding construction in the Ricci flow case by Stolarski [Sto]. More precisely, we used the Ważewski box argument to find a good perturbation using the eigenfunctions of the stability operator of the Gaussian functional. The goal is to make the mean curvature flow starting from a perturbation of a doubling of an asymptotically conical shrinker develop a singularity modeled on the chosen shrinker itself.
There are three main difficulties in the case of mean curvature flows. The first one is that we have to write the flow as a normal graph over the shrinker to use the tools from analysis, while in the Ricci flow case [Sto], one only needs to pull back tensors. The second difficulty is that we have to guarantee the embeddedness of the modified doubling of an asymptotically conical shrinker and control the curvature bound on the gluing part, while in the Ricci flow case, we can use cutoff functions to modify the metric directly. The third difficulty is that the shrinker quantity is more nonlinear than the Ricci curvature and hence complicates the arguments.
This paper is organized as follows. In Section 2, we provide preliminary notations and results for asymptotically conical self-shrinkers and talk about the evolution equations of related immersions. In Section 3, we give an explicit construction of an embedded doubling of an asymptotically conical shrinker. In Sections 4 and 5, we talk about the interior estimates and constructions of barrier functions to get the preservation and the improvement of $C^{2}$ bounds. In Section 6, we use the Ważewski box argument to finish the proof.
Acknowledgement
The authors are grateful to Prof. Toby Colding and Prof. Bill Minicozzi for their constant support and several inspiring conversations. They would also like to thank Max Stolarski for helpful conversations. During the project, Lee was partially supported by NSF Grant DMS 2005345 and Zhao was supported by NSF Grant DMS 1812142 and NSF Grant DMS 1811267.
2. Preliminaries facts and evolution equations
2.1. Stability operators on asymptotically conical shrinkers
Let $\Sigma$ be an asymptotically conical shrinker in $\mathbb{R}^{n+1}.$ In this paper, this means that its asymptotic cone
(2.1) $C=\lim\limits_{t\to 0}\sqrt{-t}\Sigma$
is a regular cone, i.e., a cone whose singular set consists of the origin. This immediately implies a curvature bound
(2.2) $\left|\nabla^{i}A\right|\leq\frac{C_{j}}{|x|^{j+1}}$
for some $C_{j}=C_{j}(\Sigma)<\infty.$ One can see [W14] for a proof of this.
The weighted Laplacian operator $\mathcal{L}$ with respect to the Gaussian functional and the stability operator of the Gaussian functional are given by (cf. [CM12])
$\mathcal{L}u$ $:=\Delta u-\frac{1}{2}\langle x^{T},\nabla u\rangle,\text{ and}$
$Lu$ $:=\mathcal{L}u+\left(|A|^{2}+\frac{1}{2}\right)u,$
where $\Delta$ is the Laplacian on $\Sigma$ and $A$ is the second fundamental form of $\Sigma.$ It is a self-adjoint operation when acting on the $W^{1,2}$ Gaussian space with respect to the weight $e^{-\frac{|x|^{2}}{4}}dV_{g}.$
The corresponding $L^{2}$ and $H^{k}$ spaces will be denoted by $L^{2}_{W}(\Sigma)$ and $H^{k}_{W}(\Sigma).$ As in the Ricci flow case ([Sto, Lemma 2.16]), the inclusion map given by $H^{1}_{W}(\Sigma)\to L^{2}_{W}(\Sigma)$ is a compact embedding. This is rigorously proven in [BW17, Appendix B] (cf. [CS, LZ]). In [CS], it is also shown that $L\colon H^{2}_{W}\left(\Sigma\right)\to L^{2}_{W}(\Sigma)$ is a bounded operator. Hence, elliptic theories implies that the spectrum of $L$ behaves as nicely as we expected.
Proposition 2.3.
There exists an $L^{2}_{W}$-orthonormal basis $\{h_{i}\}_{i\in\mathbb{N}}$ of $L^{2}_{W}(\Sigma)$ such that for all $i\in\mathbb{N},$ $h_{i}$ is an eigenfunction of $\mathcal{L}$ with eigenvalue $\lambda_{i},$ that is, $\mathcal{L}h_{i}=-\lambda_{i}h_{i}.$ The eigenvalues $\lambda_{i}$’s satisfy $\lambda_{1}<\lambda_{2}\leq\lambda_{3}\leq\cdots$ tending to +,$\infty$,+ ∞ , and each eigenspace is finite dimensional.
The next thing we need is the nice decaying property of these eigenfunctions. Similar to the eigenfunctions of the weighted Lichnerowicz operator, one expects that the $L$-eigenfunctions also decay in a suitable manner. This part follows from a general result proven by Colding-Minicozzi in [CM21]. A simpler case for $\mathcal{L}$-eigenfunctions was dealt with by Colding-Minicozzi in [CM20].
Theorem 2.4.
Let $\Sigma$ be an asymptotically conical shrinker in $\mathbb{R}^{n+1}.$ Given $\delta>0$ and $\lambda\in\mathbb{R},$ there exists $C=C(\Sigma,\delta,\lambda)<\infty$ such that the following holds. If $Lh=-\lambda h$ with $\|h\|_{L^{2}_{W}}=1,$ then
$|h(x)|\leq C|x|^{\lambda+\delta+\frac{1}{2}}.$
Proof.
The key property that plays a role here is the rapid curvature decay of $\Sigma.$ In fact, if $Lh=-\lambda h,$ we have $h\mathcal{L}h=h\left(L-\left(|A|^{2}+\frac{1}{2}\right)\right)h.$ Given a positive $\delta>0,$ by (2.2), we can find $R>0$ such that $|A|^{2}\leq\delta$ on $\Sigma\setminus B_{R}.$ Then, on $\Sigma\setminus B_{R},$ we have
$h\mathcal{L}h\geq-\lambda|h|^{2}-\left(\delta+\frac{1}{2}\right)h^{2}.$
Now we apply [CM21, Theorem 3.9] by considering $f=\frac{|x|^{2}}{4},$ $b=|x|=r,$ and $\psi=Ch^{2}$ in their work and obtain the desired estimate for $h.$ Note that the $\lambda$-eigenspace is finite-dimensional, so we can find a constant dependeing on $\lambda$ working for all $\lambda$-eigenfunctions. ∎
2.2. Evolution equation of immersions
Next, we are going to view hypersurfaces as immersions into a Euclidean space. Given a hypersurface $\Sigma\subset\mathbb{R}^{n+1}$, if we view $(\Sigma^{n},\,g)$ as a Riemannian manifold, where $g=g_{\Sigma}$ is the metric induced from the Euclidean one, there is a canonical map $F=(f_{1},f_{2},\cdots,f_{n+1})\colon\Sigma\to\mathbb{R}^{n+1}$ which corresponds to the embedding. Then we have $g(X,Y)=\langle X(F),\,Y(F)\rangle.$ When $\Sigma$ is a self-shrinker, the shrinker equation now reads
(2.5) $\Delta^{g}F=-\frac{F^{\perp}}{2}.$
We claim that the right hand side of (2.5) can be expressed in terms of $\nabla^{g}f_{i}$. In fact, in normal coordinates of $(\Sigma,g)$ around a point $x$, we have $(\nabla f_{1},\,\nabla f_{2},\,\cdots,\nabla f_{n+1})$ is an $n\times(n+1)$-matrix. From the definition, we know that the normal direction is perpendicular to the row space of this matrix. From linear algebra theories, we have that the normal vector can be expressed by
$\textbf{n}_{F}=\frac{(a_{1},\cdots,a_{n+1})}{\sqrt{\sum_{j=1}^{n+1}a_{j}^{2}}}\text{ where }a_{i}=(-1)^{i+1}\det(\nabla f_{1},\cdots,\nabla f_{i-1},\nabla f_{i+1},\cdots\nabla f_{n+1}).$
We then have $F^{\perp}=\langle F,\,\textbf{n}_{F}\rangle\textbf{n}_{F}$ can be expressed in terms of $f_{i}$’s and $\nabla f_{i}$’s.
We consider another family of immersions $F_{0}:\Sigma\times I\to\mathbb{R}^{n+1}$. It induces another family of metrics
$g_{0}(t)(X,Y)=\langle X(F_{0}(\cdot,t)),\,Y(F_{0}(\cdot,t))\rangle.$
We aim at constructing a family of diffeomorphisms $\varphi_{t}$ of $\Sigma$ and functions $h$ such that
$\frac{1}{\sqrt{1-t}}F_{0}(\varphi_{t}(x),t)=F(x)+h(x,t)\textbf{n}_{F}(x)=F_{1}(x,t),$
so $\varphi_{t}(x)$ satisfies
$(\partial_{t}F_{0})(\varphi_{t}(x),t)+dF_{0}(\partial_{t}\varphi_{t}(x))(\varphi_{t}(x),t)$
$=$ $\sqrt{1-t}\,\partial_{t}{h}(x,t)\textbf{n}_{F}(x)-\frac{1}{2\sqrt{1-t}}\left(F(x)+{h}(x,t)\textbf{n}_{F}(x)\right).$
Thus, we have
$(\partial_{t}F_{0})(\varphi_{t}(x),t)+dF_{0}(\partial_{t}\varphi_{t}(x))(\varphi_{t}(x),t)+\frac{1}{2\sqrt{1-t}}F(x)$
$=$ $(\sqrt{1-t}\,\partial_{t}{h}(x,t)-\frac{1}{2\sqrt{1-t}}{h}(x,t))\textbf{n}_{F}(x).$
As we assume that locally $\frac{1}{\sqrt{1-t}}F_{1}$ can be written as a graph over $F$, we have
$\langle\textbf{n}_{F_{0}}(\varphi_{t}(x)),\textbf{n}_{F}(x)\rangle\neq 0.$
We denote
(2.6) $\textbf{n}_{F}(x)=\langle\textbf{n}_{F}(x),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle\textbf{n}_{F_{0}}(\varphi_{t}(x))+(\textbf{n}_{F}(x)-\langle\textbf{n}_{F}(x),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle\textbf{n}_{F_{0}}(\varphi_{t}(x))),$
and we can take
$dF_{0}\left(\partial_{t}\varphi_{t}(x)\right)=$ $\frac{\langle(\partial_{t}F_{0})(\varphi_{t}(x),t)+\frac{1}{2\sqrt{1-t}}F(x),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle}{\langle\textbf{n}_{F}(x),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle}(\textbf{n}_{F}(x)-\langle\textbf{n}_{F}(x),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle\textbf{n}_{F_{0}}(\varphi_{t}(x)))$
$-((\partial_{t}F_{0})(\varphi_{t}(x),t)+\frac{1}{2\sqrt{1-t}}F(x)-\langle(\partial_{t}F_{0})(\varphi_{t}(x),t)+\frac{1}{2\sqrt{1-t}}F(x),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle\textbf{n}_{F_{0}}(\varphi_{t}(x)))$
$=$ $\frac{\langle(\partial_{t}F_{0})(\varphi_{t}(x),t)+\frac{1}{2\sqrt{1-t}}F(x),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle}{\langle\textbf{n}_{F}(x),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle}\textbf{n}_{F}(x)-((\partial_{t}F_{0})(\varphi_{t}(x),t)+\frac{1}{2\sqrt{1-t}}F(x)).$
In particular, we have
(2.7) $|dF_{0}(\partial_{t}\varphi_{t}(x))|\leq\frac{\left|(\partial_{t}F_{0})(\varphi_{t}(x),t)+\frac{1}{2\sqrt{1-t}}F(x)\right|}{|\langle\textbf{n}_{F}(x),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle|}.$
Therefore, we have
(2.8) $\partial_{t}{h}(x,t)=\frac{1}{2(1-t)}{h}(x,t)+\frac{\frac{1}{\sqrt{1-t}}\langle(\partial_{t}F_{0})(\varphi_{t}(x),t),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle+\frac{1}{2(1-t)}\langle F(x),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle}{\langle\textbf{n}_{F}(x),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle}.$
Setting $e^{\tau}=\frac{1}{1-t},$ we get
$\partial_{t}=\frac{\partial\tau}{\partial t}\partial_{\tau}=\frac{1}{1-t}\partial_{\tau},$
and thus,
$\partial_{\tau}{h}(x,t)=$ $\frac{1}{2}{h}(x,t)+\frac{{\sqrt{1-t}}\langle(\partial_{t}F_{0})(\varphi_{t}(x),t),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle+\frac{1}{2}\langle F(x),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle}{\langle\textbf{n}_{F}(x),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle}$
$=$ $\frac{\langle\Delta^{g_{1}}F_{1},\textbf{n}_{F_{1}}\rangle(x,t)}{\langle\textbf{n}_{F}(x),\textbf{n}_{F_{1}}(x)\rangle}+\frac{{\sqrt{1-t}}\langle(\partial_{t}F_{0}-\Delta^{g_{0}}F_{0})(\varphi_{t}(x),t),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle}{\langle\textbf{n}_{F}(x),\textbf{n}_{F_{1}}(x)\rangle}+\frac{\frac{1}{2}\langle F_{1}(x),\textbf{n}_{F_{1}}(x)\rangle}{\langle\textbf{n}_{F}(x),\textbf{n}_{F_{1}}(x)\rangle}.$
If we denote
$F_{1}=((f_{1})_{1},(f_{1})_{2},\cdots,(f_{1})_{n+1})\text{ and }h\,\textbf{n}_{F}=(h_{1},h_{2},\cdots,h_{n+1}),$
we have
$\textbf{n}_{F_{1}}=\frac{(\widetilde{a_{1}},\cdots,\widetilde{a_{n+1}})}{\sqrt{\sum_{j=1}^{n+1}\widetilde{a_{j}^{2}}}}\text{ where }\widetilde{a_{i}}=(-1)^{i+1}\det\left(\nabla(f_{1}+h_{1}),\cdots,\widehat{\nabla(f_{i}+h_{i})},\cdots,\nabla(f_{n+1}+h_{n+1})\right).$
In addition, we denote
$b_{i}$ $=\widetilde{a_{i}}-a_{i}$
$=(-1)^{i+1}\left(\det\left(\nabla(f_{1}+h_{1}),\cdots,\widehat{\nabla(f_{i}+h_{i})},\cdots\nabla(f_{n+1}+h_{n+1})\right)-\det\left(\nabla f_{1},\cdots,\widehat{\nabla f_{i}},\cdots\nabla f_{n+1}\right)\right).$
Then, we derive
$\widetilde{\textbf{n}}:=\textbf{n}_{F_{1}}-\textbf{n}_{F}=\left(\frac{a_{i}\left(\sum_{j=1}^{n+1}-2a_{j}b_{j}-b_{j}^{2}\right)+b_{i}\sqrt{\sum_{j=1}^{n+1}a_{j}^{2}}\left(\sqrt{\sum_{j=1}^{n+1}(a_{j}+b_{j})^{2}}+\sqrt{\sum_{j=1}^{n+1}a_{j}^{2}}\right)}{\sqrt{\sum_{j=1}^{n+1}(a_{j}+b_{j})^{2}}\sqrt{\sum_{j=1}^{n+1}a_{j}^{2}}\left(\sqrt{\sum_{j=1}^{n+1}(a_{j}+b_{j})^{2}}+\sqrt{\sum_{j=1}^{n+1}a_{j}^{2}}\right)}\right)_{i=1,2,\cdots,n+1}.$
In local coordinates, we have
$\frac{\left\langle\Delta^{g_{1}}F_{1}+\frac{F_{1}^{\perp}}{2},\textbf{n}_{F_{1}}\right\rangle}{\langle\textbf{n}_{F},\textbf{n}_{F_{1}}\rangle}\textbf{n}_{F}=\frac{\left\langle\Delta^{g_{1}}F_{1}+\frac{F_{1}^{\perp}}{2},\textbf{n}_{F}\right\rangle}{\langle\textbf{n}_{F},\textbf{n}_{F_{1}}\rangle^{2}}\textbf{n}_{F}$
$=$ $\left\langle\Delta^{g}(h)+\frac{F_{1}^{\perp}-F^{\perp}}{2}+(\Delta^{g_{1}}-\Delta^{g})F_{1},\textbf{n}_{F}\right\rangle\frac{\textbf{n}_{F}}{\langle\textbf{n}_{F},\textbf{n}_{F_{1}}\rangle^{2}}$
$=$ $\left\langle\Delta^{g}(h\,\textbf{n}_{F})-\frac{1}{2}\left(\langle F,\textbf{n}_{F}\rangle\textbf{n}_{F}-\langle F_{1},\textbf{n}_{F_{1}}\rangle\textbf{n}_{F_{1}}\right)+\left(g_{1}^{ab}-g^{ab}\right)\partial_{a}\partial_{b}F_{1}-\left(g_{1}^{ab}\Gamma^{g_{1},c}_{ab}-g^{ab}\Gamma^{g,c}_{ab}\right)\partial_{c}F_{1},\textbf{n}_{F}\right\rangle\frac{\textbf{n}_{F}}{\langle\textbf{n}_{F},\textbf{n}_{F_{1}}\rangle^{2}}$
$=$ $\left\langle\Delta^{g}(h\,\textbf{n}_{F})+\frac{1}{2}\langle(h\,\textbf{n}_{F}),\textbf{n}_{F}\rangle\textbf{n}_{F}-\frac{1}{2}(\langle F_{1},\textbf{n}_{F}\rangle\textbf{n}_{F}-\langle F_{1},\textbf{n}_{F}+\widetilde{\textbf{n}}\rangle(\textbf{n}_{F}+\widetilde{\textbf{n}}))+\widetilde{(h\textbf{n}_{F})}^{ab}\partial_{a}\partial_{b}F_{1}\right.$
$\left.-\widetilde{(h\,\textbf{n}_{F})}^{ab}\Gamma^{g,c}_{ab}\partial_{c}F_{1}-g_{1}^{ab}(\Gamma^{g_{1},c}_{ab}-\Gamma^{g,c}_{ab})\partial_{c}F_{1},\textbf{n}_{F}\right\rangle\frac{\textbf{n}_{F}}{\langle\textbf{n}_{F},\textbf{n}_{F_{1}}\rangle^{2}}$
$=$ $\left\langle\Delta^{g}(h\,\textbf{n}_{F})+\frac{1}{2}\langle(h\,\textbf{n}_{F}),\textbf{n}_{F}\rangle\textbf{n}_{F}+\frac{1}{2}(\langle F+(h\,\textbf{n}_{F}),\widetilde{\textbf{n}}\rangle(\textbf{n}_{F}+\widetilde{\textbf{n}})+\langle F+(h\,\textbf{n}_{F}),\textbf{n}_{F}\rangle\widetilde{\textbf{n}})+\tilde{h}^{ab}\partial_{a}\partial_{b}F_{1}\right.$
$-\tilde{h}^{ab}\Gamma^{g,c}_{ab}\partial_{c}F_{1}-\frac{1}{2}((\tilde{h}^{ab}+g^{ab})(g^{cd}+\tilde{h}^{cd})(\hat{h}_{da,b}+\hat{h}_{db,a}-\hat{h}_{ab,d})-\tilde{h}^{ab}\tilde{h}^{cd}(g_{da,b}+g_{db,a}-g_{ab,d}))\partial_{c}F_{1}$
$\left.-\frac{1}{2}(g^{ab}\tilde{h}^{cd}(g_{da,b}+g_{db,a}-g_{ab,d}))\partial_{c}F_{1},\textbf{n}_{F}\right\rangle\frac{\textbf{n}_{F}}{\langle\textbf{n}_{F},\textbf{n}_{F_{1}}\rangle^{2}}.$
Thus, we have
$\frac{\partial(h\,\textbf{n}_{F})}{\partial\tau}$
$=$ $\left\langle\Delta^{g}(h\,\textbf{n}_{F})+\frac{1}{2}h\,\textbf{n}_{F}+\frac{1}{2}(\langle F,\widetilde{\textbf{n}}\rangle\textbf{n}_{F}+\langle F,\textbf{n}_{F}\rangle\widetilde{\textbf{n}})+\tilde{h}^{ab}\nabla^{g}_{a}\nabla^{g}_{b}F\right.$
$-\frac{1}{2}\left(g^{ab}g^{cd}\left(\hat{h}_{da,b}+\hat{h}_{db,a}-\hat{h}_{ab,d}\right)+g^{ab}\tilde{h}^{cd}(g_{da,b}+g_{db,a}-g_{ab,d})\right)\partial_{c}F+\frac{1}{2}(\langle h\,\textbf{n}_{F},\widetilde{\textbf{n}}\rangle(\textbf{n}_{F}+\widetilde{\textbf{n}})+\langle F,\widetilde{\textbf{n}}\rangle\widetilde{\textbf{n}}+h\widetilde{\textbf{n}})$
$+\tilde{h}^{ab}\nabla^{g}_{a}\nabla^{g}_{b}(h\,\textbf{n}_{F})-\frac{1}{2}\left(g^{ab}g^{cd}\left(\hat{h}_{da,b}+\hat{h}_{db,a}-\hat{h}_{ab,d}\right)+g^{ab}\tilde{h}^{cd}\left(g_{da,b}+g_{db,a}-g_{ab,d}\right)\right)\partial_{c}(h\,\textbf{n}_{F})$
$-\left.\frac{1}{2}((\tilde{h}^{ab}g^{cd}+g^{ab}\tilde{h}^{cd})(\hat{h}_{da,b}+\hat{h}_{db,a}-\hat{h}_{ab,d})-\tilde{h}^{ab}\tilde{h}^{cd}(g_{da,b}+g_{db,a}-g_{ab,d}))(\partial_{c}F+\partial_{c}(h\,\textbf{n}_{F}))),\textbf{n}_{F}\right\rangle\frac{\textbf{n}_{F}}{\langle\textbf{n}_{F},\textbf{n}_{F_{1}}\rangle^{2}}$
$+\frac{{\sqrt{1-t}}\langle(\partial_{t}F_{0}-\Delta^{g_{0}}F_{0})(\varphi_{t}(x),t),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle}{\langle\textbf{n}_{F}(x),\textbf{n}_{F_{1}}(x)\rangle}\textbf{n}_{F}$
$=$ $\left\langle\Delta^{g}h\cdot\textbf{n}_{F}-h|A|^{2}\textbf{n}_{F}+\frac{1}{2}h\,\textbf{n}_{F}+\frac{1}{2}(\langle F,\widetilde{\textbf{n}}\rangle\textbf{n}_{F}+\langle F,\textbf{n}_{F}\rangle\widetilde{\textbf{n}})+\tilde{h}^{ab}\nabla^{g}_{a}\nabla^{g}_{b}F\right.$
$+\frac{1}{2}(\langle h\,\textbf{n}_{F},\widetilde{\textbf{n}}\rangle(\textbf{n}_{F}+\widetilde{\textbf{n}})+\langle F,\widetilde{\textbf{n}}\rangle\widetilde{\textbf{n}}+h\widetilde{\textbf{n}})+\tilde{h}^{ab}\nabla^{g}_{a}\nabla^{g}_{b}(h\,\textbf{n}_{F})$
$-\frac{1}{2}\left(g^{ab}g^{cd}(\hat{h}_{da,b}+\hat{h}_{db,a}-\hat{h}_{ab,d})+g^{ab}\tilde{h}^{cd}(g_{da,b}+g_{db,a}-g_{ab,d})\right)\partial_{c}h\cdot\textbf{n}_{F}$
$-\left.\frac{1}{2}\left((\tilde{h}^{ab}g^{cd}+g^{ab}\tilde{h}^{cd})(\hat{h}_{da,b}+\hat{h}_{db,a}-\hat{h}_{ab,d})-\tilde{h}^{ab}\tilde{h}^{cd}(g_{da,b}+g_{db,a}-g_{ab,d})\right)\partial_{c}h\cdot\textbf{n}_{F},\textbf{n}_{F}\right\rangle\frac{\textbf{n}_{F}}{(1+\langle\textbf{n}_{F},\widetilde{\textbf{n}}\rangle)^{2}}$
$+\frac{{\sqrt{1-t}}\langle(\partial_{t}F_{0}-\Delta^{g_{0}}F_{0})(\varphi_{t}(x),t),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle}{\langle\textbf{n}_{F}(x),\textbf{n}_{F_{1}}(x)\rangle}\textbf{n}_{F},$
where $\tilde{h}^{ab}=g^{ab}_{1}-g^{ab}$, $\hat{h}_{ab}=(g_{1})_{ab}-g_{ab}$.
As in the beginning, we denote the second fundamental form on the image of $F$ as $A_{ij}$. If we work in normal coordinates around a point, we have
(2.9) $\widetilde{\textbf{n}}$ $=-\nabla^{g}{h}+C_{1}(F,\nabla F,\nabla^{2}F,h,\nabla h)*\nabla(h\,\textbf{n}_{F})*\nabla(h\,\textbf{n}_{F}),$
$\hat{h}_{ab}$ $=\langle\partial_{a}F,\partial_{b}(h\,\textbf{n}_{F})\rangle+\langle\partial_{a}(h\,\textbf{n}_{F}),\partial_{b}F\rangle+\langle\partial_{a}(h\,\textbf{n}_{F}),\partial_{b}(h\,\textbf{n}_{F})\rangle=2hA_{ab}+\langle\partial_{a}h,\partial_{b}h\rangle+h^{2}A_{ia}A_{bj}g^{ij},$
$\tilde{h}^{ab}$ $=-2hA^{ab}+C_{2,ab}(F,\nabla F,\nabla^{2}F,h,\nabla h)*\nabla h*\nabla h+C_{3,ab}(F,\nabla F,\nabla^{2}F,h,\nabla h)*hA,$
where
$\|C_{1}(F,\nabla F,\nabla^{2}F,h,\nabla h)\|_{C^{0}}+\sum_{a,b}\|C_{2,ab}(F,\nabla F,\nabla^{2}F,h,\nabla h)\|_{C^{0}}+\|C_{3,ab}(F,\nabla F,\nabla^{2}F,h,\nabla h)\|_{C^{0}}\leq C_{0}\|h\|_{C^{1}}.$
Also, note that $\Delta^{g}\textbf{n}_{F}=-|A|^{2}\textbf{n}_{F}+\nabla^{g}H$ and $A_{ij,k}=A_{ik,j}$ in $\mathbb{R}^{n+1}$. Here, we use the notation $\textbf{H}=\Delta^{g}F$ and $H=-\textbf{H},\textbf{n}_{F}\rangle=g^{ij}A_{ij},$ so in normal coordinates, we have
$\left\langle\Delta^{g}h\cdot\textbf{n}_{F}-h|A|^{2}\textbf{n}_{F}+\frac{1}{2}h\textbf{n}_{F}+\frac{1}{2}(\langle F,\widetilde{\textbf{n}}\rangle\textbf{n}_{F}+\langle F,\textbf{n}_{F}\rangle\widetilde{\textbf{n}})+\tilde{h}^{ab}\nabla^{g}_{a}\nabla^{g}_{b}F,\textbf{n}_{F}\right\rangle$
$=$ $\left\langle(\Delta^{g}h-|A|^{2}h)\textbf{n}_{F}+\frac{1}{2}h\,\textbf{n}_{F}+2h|A|^{2}\textbf{n}_{F}+\frac{1}{2}(\langle F,-\nabla^{g}{h}\rangle\textbf{n}_{F}+\langle F,\textbf{n}_{F}\rangle(-\nabla^{g}{h}))\right.$
$\left.+C_{1}^{\prime}(F,\nabla F,h,\nabla h)*\nabla(h\,\textbf{n}_{F})+C_{2}^{\prime}(F,\nabla F,h,\nabla h)h\,\textbf{n}_{F},\textbf{n}_{F}\right\rangle$
$=$ $\left(\Delta^{g}h+|A|^{2}h-\frac{1}{2}\langle F,\nabla^{g}h\rangle+\frac{1}{2}h\right)+C^{\prime}_{1}(F,\nabla F,h,\nabla h)*\nabla(h\,\textbf{n}_{F})+C_{2}^{\prime}(F,\nabla F,h,\nabla h)h.$
We recall the stability operator $L$ for the Gaussian function mentioned before, written using the embedding $F,$
$Lf=\Delta^{g}f+|A|^{2}f-\frac{1}{2}\langle F,\nabla^{g}f\rangle+\frac{1}{2}f.$
Combining all the calculations above, we can write the evolution equation of $h$ as
(2.10) $\frac{\partial h}{\partial\tau}=$ $Lh+\frac{{\sqrt{1-t}}\langle(\partial_{t}F_{0}-\Delta^{g_{0}}F_{0})(\varphi_{t}(x),t),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle}{\langle\textbf{n}_{F}(x),\textbf{n}_{F_{1}}(x)\rangle}$
$+C_{3}(F,\nabla F,h,\nabla h)*h+C_{4}(F,\nabla F,h,\nabla h)*\nabla h+C_{5}(F,\nabla F,h,\nabla h)*\nabla^{2}h$
$=$ $:Lh+\mathcal{Q}_{1}+\mathcal{Q}_{2},$
where
$\mathcal{Q}_{1}=$ ${\sqrt{1-t}}\left\langle(\partial_{t}F_{0}-\Delta^{g_{0}}F_{0})(\varphi_{t}(x),t),\textbf{n}_{F_{0}}(\varphi_{t}(x))\right\rangle,\text{ and}$
$\mathcal{Q}_{2}=$ $C_{3}(F,\nabla F,h,\nabla h)*h+C_{4}(F,\nabla F,h,\nabla h)*\nabla h+C_{5}(F,\nabla F,h,\nabla h)*\nabla^{2}h$
$+{\sqrt{1-t}}\left\langle(\partial_{t}F_{0}-\Delta^{g_{0}}F_{0})(\varphi_{t}(x),t),\textbf{n}_{F_{0}}(\varphi_{t}(x))\right\rangle\left(\frac{1}{{\langle\textbf{n}_{F}(x),\textbf{n}_{F_{1}}(x)\rangle}}-1\right).$
We remark that in normal coordinates, the second order term in $\mathcal{Q}_{2}$ corresponds to
(2.11) $(\tilde{h}^{ab}\partial_{a}\partial_{b}h-g^{ab}g^{cd}\partial_{a}\partial_{b}h\,\partial_{d}h\,\partial_{c}h-(\tilde{h}^{ab}g^{cd}+g^{ab}\tilde{h}^{cd})\partial_{a}\partial_{b}h\,\partial_{d}h\,\partial_{c}h)\frac{\textbf{n}_{F}}{(1+\langle\textbf{n}_{F},\widetilde{\textbf{n}}\rangle)^{2}}$
$=$ $\frac{\tilde{h}^{ab}(1-|\nabla h|^{2})-(|\nabla h|^{2}+\tilde{h}^{cd}\partial_{c}h\,\partial_{d}h)g^{ab}}{(1+\langle\textbf{n}_{F},\widetilde{\textbf{n}}\rangle)^{2}}\partial_{a}\partial_{b}h\cdot\textbf{n}_{F}.$
3. Constructions of doubling of asymptotically conical shrinkers
We are going to construct a closed hypersurface which is a double of a large compact set of the concerned asymptotically conical shrinker $\Sigma$. The main idea is similar to [Sto], but here the process will be more technical since we are working on extrinsic hypersurfaces.
As in (2.1), denoting the asymptotic cone of $\Sigma$ by $C$, from [W14, CS], we know that outside a large compact set, we can parameterize $\Sigma\setminus B_{R}(0)$ as
$F_{0}(r,\theta)=r\theta+f(r,\theta)\textbf{n}$
over $C$, where $\theta$ is the parametrization for the link of $C$ and n is the normal vector of $C$ at $(r,\theta)$.
We choose a smooth cutoff function $\eta\colon\mathbb{R}\to\mathbb{R}$ such that $\eta=\left\{\begin{array}[]{cc}1,\,\,x<0\\ 0,\,\,x>1\end{array}\right..$ Take $F_{1}(r,\theta)=r\theta+\eta(\frac{r-R}{R})f(r,\theta)\textbf{n}$. Note that to glue two copies, we need to modify the end such that it becomes cylindrical.
The first step is to move the link to a hemisphere. As the link of $C$ is a codimension one submanifold of $S^{n-1}$, we can pick a vector $\mathbf{v}\in S^{n-1}$ such that $-\mathbf{v}$ is not contained in the link of $C$. Initially, we have $F_{1}(r,\theta)=r\theta$ for $r>2R.$ We take
$F_{2}(r,\theta)=r\cdot\frac{\left(\frac{2}{3}\eta\left(\frac{r-2R}{R}\right)+\frac{1}{3}\right)\theta+\left(\frac{2}{3}-\frac{2}{3}\eta\left(\frac{r-2R}{R}\right)\right)\mathbf{v}}{\left|\left(\frac{2}{3}\eta\left(\frac{r-2R}{R}\right)+\frac{1}{3}\right)\theta+\left(\frac{2}{3}-\frac{2}{3}\eta\left(\frac{r-2R}{R}\right)\right)\mathbf{v}\right|}$
so that for $r\geq 3R,$ the image is contained in a hemisphere. Consequently, we can define
$F_{3}(r,\theta)=\begin{cases}r\theta+\eta(\frac{r-R}{R})f(r,\theta)\textbf{n}&\text{if }r\leq 2R\\ r\cdot\frac{\left(\frac{2}{3}\eta\left(\frac{r-2R}{R}\right)+\frac{1}{3}\right)\theta+\left(\frac{2}{3}-\frac{2}{3}\eta\left(\frac{r-2R}{R}\right)\right)\mathbf{v}}{\left|\left(\frac{2}{3}\eta\left(\frac{r-2R}{R}\right)+\frac{1}{3}\right)\theta+\left(\frac{2}{3}-\frac{2}{3}\eta\left(\frac{r-2R}{R}\right)\right)\mathbf{v}\right|}&\text{if }r\geq 2R\end{cases}.$
The next step is to modify the end such that the radial derivative is in the same direction. To this end, we take
[…]
The last step is to make the end cylindrical and end at the same plane so that we can do a reflection to double it. Denote $B:=\sup_{\theta}\langle F_{4}(4R,\theta),\,\mathbf{v}\rangle,$ and consider
[…]
From the constructions, the $F_{3}$ part lies in a hemisphere and controls the distance to origin. Thus, we know that $F$ is globally an embedding, and after reflecting it with respect to the hyperplane $\{x\in\mathbb{R}^{n+1}\,|\,\langle x,\mathbf{v}\rangle=B+R\},$ we get a closed embedded hypersurface.
4. Interior estimates
We are going to establish some interior estimates, as in the Ricci flow case developed in [Sto]. To get the required Hölder estimates in the case of mean curvature flows, we will modify the estimates in [Bam, Proposition 2.5], [App, Lemma 4.4], and [Sto, Proposition C.4.]. For completeness, we sketch their proofs here.
[…]
Take $02r\geq s>$ 0 and consider $\Omega=B_{r}\times[0,s^{2}]$ and $\Omega^{\prime}=B_{2r}\times[0,s^{2}]$. Assume $u\in C^{2}(\Omega^{\prime})$ satisfies
$(\partial_{t}-L)u=Q[u]+F=$ $r^{-2}f_{1}(r^{-1}x,r^{-2}t,u,r\nabla u)\cdot u+r^{-1}f_{2}(r^{-1}x,r^{-2}t,u,r\nabla u)\cdot\nabla u$
$+f_{3}(r^{-1}x,r^{-2}t,u,r\nabla u)\cdot\nabla u\otimes\nabla u+f_{4}(r^{-1}x,r^{-2}t,u,r\nabla u)\cdot u\otimes\nabla^{2}u+F(x,t)$
with $u(\cdot,0)=u_{0}$, where $f_{1},f_{2},f_{3},$ and $f_{4}$ are smooth. The operator $L$ is in the form
$Lu=a^{ij}(x,t)\partial^{2}_{ij}u+b^{i}(x,t)\partial_{i}u+c(x,t)u$
and the coefficients satisfy the following bounds, for $m\geq 1,$
$\frac{1}{\Lambda}\leq a^{ij}\leq\Lambda,\,\|a^{ij}\|_{C^{2m-2,\alpha}(\Omega^{\prime})}\leq\Lambda,\,\|b^{i}\|_{C^{2m-2,\alpha}(\Omega^{\prime})}\leq r^{-1}\Lambda,\,\text{ and }\|c\|_{C^{2m-2,\alpha}(\Omega^{\prime})}\leq r^{-2}\Lambda.$
Then there exist constants $\varepsilon>0$ and $C<\infty$ depending on $n,m,\alpha,\Lambda,$ and $f_{i}$ such that if
$\|u\|_{c^{0}(\Omega^{\prime})}+\|u_{0}\|_{C^{2m,\alpha}(B_{2r})}+r^{2}\|F\|_{C^{2m-2,\alpha}(\Omega^{\prime})}\leq\varepsilon,$
then
$\|u\|_{C^{2m,\alpha}(\Omega)}\leq C\left(\|u\|_{C^{0}(\Omega^{\prime})}+\|u_{0}\|_{C^{2m,\alpha}(B_{2r})}+r^{2}\|F\|_{C^{2m-2,\alpha}(\Omega^{\prime})}\right).$
Proof.
As in [Sto], we define the weighted norm as
$\|u\|^{(\theta)}_{C^{2m,\alpha}(\Omega)}:=\sum_{i+2j\leq 2m}(r_{\Omega}\theta)^{i+2j}(\|\nabla^{i}\partial^{j}_{t}u\|_{C^{0}(\Omega)}+(r_{\Omega}\theta)^{\alpha}[\nabla^{i}\partial^{j}_{t}u]_{\alpha,\frac{\alpha}{2}}).$
Then, from the standard parabolic estimates (cf. [Kry]), we have
$\|u\|_{C^{2m,\alpha}(B_{r}\times[0,s^{2}])}\leq C\left((r\theta)^{2}\|Q[u]+F\|_{C^{2m-2,\alpha}(B_{(1+\theta)r}\times[0,s^{2}])}^{(\theta)}+\|u\|^{(\theta)}_{C^{0}(B_{(1+\theta)r}\times[0,s^{2}])}+\|u_{0}\|^{(\theta)}_{C^{2m,\alpha}(B_{(1+\theta)r})}\right).$
After rescaling, we may assume $r=1$. Define
$H$ $:=\|u\|_{C^{0}(B_{2}\times[0,s^{2}])}+\|u_{0}\|_{C^{2m,\alpha}(B_{2})}+\|F\|_{C^{2m-2,\alpha}(B_{2}\times[0,s^{2}])},$
$r_{k}$ $=2-2^{-k},\,\theta_{k}=\frac{r_{k+1}-r_{k}}{r_{k}},\,\Omega=B_{r_{k}}\times[0,s^{2}],\text{ and}$
$a_{k}$ $:=\|u\|^{(\theta_{k})}_{C^{2m,\alpha}(\Omega_{k})}\leq C\left(\theta_{k}^{2}\|Q[u]\|^{(\theta_{k+1})}_{C^{2m-2,\alpha}(\Omega_{k+1})}+H\right).$
For $i=1,\cdots,4,$ we have
$\|f_{i}(x,t,u,r\nabla u)\|_{C^{2m-2,\alpha}(\Omega_{k+1})}^{(\theta_{k+1})}\leq C\left(1+(\|u\|_{C^{2m,\alpha}(\Omega_{k+1})}^{(\theta_{k+1})})^{2m-1}\right)\leq C\left(1+a_{k+1}^{2m-1}\right),$
so by plugging in the estimates, we get
$a_{k}\leq C\left(\theta_{k+1}a_{k+1}+a_{k+1}^{2}+a_{k+1}^{2m}+a_{k+1}^{2m+1}+H\right).$
Take $k_{0}$ such that for $k\geq k_{0}$, $\theta_{k+1}\leq\frac{1}{16C_{0}^{2}}$. Then for $b_{k}:=\frac{a_{k}}{H}$, $K\geq k_{0},$ and $H\leq\varepsilon,$ we have
$b_{k}\leq\frac{1}{16}b_{k+1}+\frac{1}{16C_{0}}b_{k+1}^{2}+\frac{1}{16^{2m-1}C_{0}^{4m-3}}b_{k+1}^{2m}+\frac{1}{16^{2m}C_{0}^{4m-1}}b_{k+1}^{2m+1}+C_{0}.$
As $\lim_{k\to\infty}b_{k}\leq 1<2C_{0}$, by induction, we know that $b_{k_{0}}<2C_{0}$. Hence $a_{0}\leq C^{\prime}H$, which completes the proof. ∎
From Proposition 4, we have the following estimate.
[…]
Given $h$ satisfying (2.10) on $\Omega^{\prime}\times(\tau_{0},\tau_{2})$, let $\Omega\Subset\Omega^{\prime}\Subset\Sigma$, $m\in\mathbb{N}$, $\tau_{0}\leq\tau_{1}<\tau_{2},$ and $\mathcal{Q}_{1}=0$ on $\Omega^{\prime}\times[\tau_{0},\tau_{2}]$. Then there exist constants $\varepsilon=\varepsilon(n,m)>0$ and $C=C\left(n,\Sigma,F,{\text{dist}}(\Omega,M\setminus\Omega^{\prime}),\tau_{1}-\tau_{0}\right)<\infty$ with the following estimates.
(1) If $\tau_{1}>\tau_{0}$, then
$\|h\|_{L^{\infty}(\Omega^{\prime}\times[\tau_{0},\tau_{2}])}<\varepsilon\Rightarrow\|h\|_{C^{m}(\Omega\times[\tau_{1},\tau_{2}])}\leq C\|h\|_{L^{\infty}(\Omega^{\prime}\times[\tau_{0},\tau_{2}])}.$
(2) If $\tau_{1}=\tau_{0}$, then
$\|h\|_{L^{\infty}(\Omega^{\prime}\times[\tau_{0},\tau_{2}])}+\|h\|_{C^{m+1}(\Omega^{\prime}\times\{\tau_{0}\})}<\varepsilon\Rightarrow\|h\|_{C^{m}(\Omega\times[\tau_{0},\tau_{2}])}\leq C\left(\|h\|_{L^{\infty}(\Omega^{\prime}\times[\tau_{0},\tau_{2}])}+\|h\|_{C^{m+1}(\Omega^{\prime}\times\{\tau_{0}\})}\right).$
On the other hand, we also have interior integral estimates. Recall that we use $L^{2}_{W}$ and $H^{k}_{W}$ to be the weighted $L^{2}$ and $H^{k}$ spaces with the weight given by $e^{-\frac{|F|^{2}}{4}}.$
[…]
Given $h$ satisfying (2.10) on $\Omega^{\prime}\times(\tau_{0},\tau_{2})$, let $\Omega\Subset\Omega^{\prime}\Subset\Sigma$, $m\in\mathbb{N}$, $\tau_{0}\leq\tau_{1}<\tau_{2},$ and $\mathcal{Q}_{1}=0$ on $\Omega^{\prime}\times[\tau_{0},\tau_{2}]$. Then there exist constants $\varepsilon=\varepsilon(n,m)>0$ and $C=C(n,\Sigma,F,{\text{dist}}(\Omega,M\setminus\Omega^{\prime}),\tau_{1}-\tau_{0})<\infty$ with the following estimates.
(1) If $\tau_{1}>\tau_{0}$, then
$\|h\|_{C^{m+2}(\Omega^{\prime}\times[\tau_{0},\tau_{2}])}<\varepsilon\Rightarrow\sup_{\tau\in[\tau_{1},\tau_{2}]}\|\nabla^{m}h(\tau)\|_{L^{2}_{W}(\Omega)}+\|h\|_{H_{W}^{m+1}(\Omega\times[\tau_{1},\tau_{2}])}\leq C\|h\|_{H_{W}^{m}(\Omega^{\prime}\times[\tau_{0},\tau_{2}])}.$
(2) If $\tau_{1}=\tau_{0}$, then
$\|h\|_{C^{m+2}(\Omega^{\prime}\times[\tau_{0},\tau_{2}])}<\varepsilon$
$\Rightarrow$ $\sup_{\tau\in[\tau_{0},\tau_{2}]}\|\nabla^{m}h(\tau)\|_{L^{2}_{W}(\Omega)}+\|h\|_{H_{W}^{m+1}(\Omega\times[\tau_{0},\tau_{2}])}\leq C\left(\|h(\tau_{0})\|_{H_{W}^{m}(\Omega^{\prime})}+\|h\|_{H_{W}^{m}(\Omega^{\prime}\times[\tau_{0},\tau_{2}])}\right).$
Proof.
The proof is essentially the same as [Sto, Lemma 6.3]. The only difference is that we need to assume that one more order derivative is bounded as the coefficient of (6.2) is more nonlinear.
From (2.10), (6.2), and the assumption $\mathcal{Q}_{1}=0$ on $\Omega^{\prime}\times[\tau_{0},\tau_{2}]$, we have
$\partial_{t}h=\mathcal{L}h+\sum_{l=1}^{1}B_{l}^{0}*\nabla^{l}h-{\rm div}_{\frac{|F|^{2}}{4}}\left(\hat{h}^{0}*\nabla h\right),$
where $\mathcal{L}=\Delta_{\frac{|F|^{2}}{4}}$ is the drift Laplacian and there exists $C_{0},$ and for any $\delta>0$, there exists $\varepsilon>0$ such that
$\|h(\tau)\|_{C^{2}(\Omega^{\prime})}<\varepsilon\Rightarrow\|B_{1}^{0}\|_{L^{\infty}(\Omega^{\prime})}+\|B_{0}^{0}\|_{L^{\infty}(\Omega^{\prime})}<C_{0},\,\|\hat{h}^{0}\|_{L^{\infty}(\Omega^{\prime})}\leq\delta.$
By induction, we have
$\partial_{t}\nabla^{m}h=\mathcal{L}\nabla^{m}h+\sum_{l=1}^{m+1}B_{l}^{m}*\nabla^{l}h-{\rm div}_{\frac{|F|^{2}}{4}}\left(\hat{h}^{m}*\nabla^{m+1}h\right),$
where there exists $C_{0}$ and for any $\delta>0$, there exists $\varepsilon>0$ such that
$\|h(\tau)\|_{C^{m+2}(\Omega^{\prime})}<\varepsilon\Rightarrow\sum_{l=0}^{m+1}\|B_{l}^{m}\|_{L^{\infty}(\Omega^{\prime})}<C_{m},\,\|\hat{h}^{m}\|_{L^{\infty}(\Omega^{\prime})}\leq\delta.$
Then the rest of the proof follows from integration by parts and Young’s inequality, cf. [Sto, Lemma 6.3]. ∎
Finally, we arrive at the local interior estimate we will use later.
[…]
Given $\Gamma>0$ and $h$ satisfying (2.10) on $\Sigma\times(0,\tau_{1})$, assume that $\mathcal{Q}_{1}=0$ on $\left\{\frac{|F|^{2}}{4}<2\Gamma e^{\tau_{0}}\right\}\times[0,\tau_{1}]$ and
$\|h\|_{L^{\infty}(\Sigma\times[0,\tau_{1}])}$ $<\varepsilon,$
$\|h(\tau)\|_{L_{\frac{|f|^{2}}{4}}^{2}(\Sigma)}$ $\leq\mu e^{\lambda_{*}(\tau+\tau_{0})}\,\,\text{ for all}\,\,\tau\in[0,\tau_{1}],\text{ and}$
$\sum_{l=0}^{4n+3}\left|\nabla^{l}h\right|_{g}(x,0)$ $\leq C_{0}\bar{p}e^{\lambda_{*}\tau_{0}}\,\,\text{ for all}\,\,x\in\left\{\frac{|F|^{2}}{4}<2\Gamma e^{\tau_{0}}\right\}\text{ and }\tau=0.$
Then for $\Gamma>\Gamma_{0}(n,\Sigma,F)$, $\varepsilon<\varepsilon_{0}(n,\Sigma,F,\Gamma),$ and $\tau_{0}>C(n,\Sigma,F,\lambda_{*},\Gamma,C_{0}\bar{p})$, there exists $C_{1}=C_{1}(n,\Sigma,F,\lambda_{*},\Gamma)<\infty$ such that
$\sum_{l=0}^{2}|\nabla^{g}h|_{g}(x,\tau)\leq C_{1}(\mu+C_{0}\bar{p})e^{\lambda_{*}(\tau+\tau_{0})}\,\,\text{ for all}\,\,(x,\tau)\in\left\{\frac{|F|^{2}}{4}<\Gamma e^{\tau_{0}}\right\}\times[0,\tau_{1}].$
Note that we can also use Proposition 4 to get large scale interior estimates by rescaling. More precisely, consider a family of diffeomorphisms $\varphi_{s}\colon M\to M$ given by
$\partial_{s}\varphi_{s}=\nabla^{g}\left(\frac{|F|^{2}}{4}\right)\circ\varphi_{s}$
and define $\mathcal{H}(t)=\varphi_{t}^{*}h(t).$ The evolution equation of $\mathcal{H}$ is then
(4.1) $\frac{\partial\mathcal{H}}{\partial t}(s)=$ $\left(\Delta^{g}+|A|^{2}+\frac{1}{2}\right)\mathcal{H}(s)+\varphi_{s}^{*}\mathcal{Q}_{1}+\varphi_{s}^{*}\mathcal{Q}_{2}.$
This helps us to get rid of the unbounded coefficient term $-\frac{1}{2}\langle F,\nabla^{g}f\rangle$ in $L$. Also, note that $0\leq\left|\nabla^{g}(\frac{|F|^{2}}{4})\right|^{2}\leq\frac{|F|^{2}}{4}$ gives us that for $s_{1}<s_{2},$
$\frac{|F|^{2}}{4}(\varphi_{s_{1}}(x))\leq\frac{|F|^{2}}{4}(\varphi_{s_{2}}(x))$
and
$e^{-s_{1}}\frac{|F|^{2}}{4}(\varphi_{s_{1}}(x))\geq e^{-s_{2}}\frac{|F|^{2}}{4}(\varphi_{s_{2}}(x)).$
Thus, Proposition 4 implies the large scale interior estimates.
[…]
Given $h$ satisfying (2.10) and $\mathcal{Q}_{1}=0$ on $\left\{(x,\tau)\in\Sigma\times(0,\tau_{1}):\,\Gamma e^{\tau_{0}}<\frac{|F|^{2}}{4}<\gamma e^{\tau+\tau_{0}}\right\},$ assume that
$|h|_{g}$ $\leq C_{0}e^{-k(\tau+\tau_{0})}\left(\frac{|F|^{2}}{4}(x)\right)^{k}\,\,\text{on}\,\,\left\{(x,\tau)\in\Sigma\times(0,\tau_{1}):\,\Gamma e^{\tau_{0}}<\frac{|F|^{2}}{4}<\gamma e^{\tau+\tau_{0}}\right\},\text{ and}$
$\sum_{l=0}^{3}|\nabla^{l}h|_{g}(x,\tau_{0})$ $\leq C_{0}e^{-k\tau_{0}}\left(\frac{|F|^{2}}{4}(x)\right)^{k}\,\,\text{for all}\,\,x\in\left\{\Gamma e^{\tau_{0}}<\frac{|F|^{2}}{4}<\gamma e^{\tau+\tau_{0}}\right\}.$
Then for $\Gamma>\Gamma_{0}(n,\Sigma,F)$ and $C_{0}\gamma^{k}<C(n,\Sigma,F)$, there exists $C_{1}=C_{1}(n,\Sigma,F,k)<\infty$ such that
$\sum_{l=0}^{2}|\nabla^{g}h|_{g}(x,\tau)\leq C_{1}C_{0}e^{-k(\tau+\tau_{0})}\left(\frac{|F|^{2}}{4}(x)\right)^{k}\,\,\text{on}\,\,\left\{(x,\tau)\in\Sigma\times[0,\tau_{1}]:\,4\Gamma e^{\tau_{0}}\leq\frac{|F|^{2}}{4}(x)\leq\frac{1}{4}\gamma e^{\tau+\tau_{0}}\right\}.$
Similarly, on the outer region, we have the following estimate.
[…]
Given $h$ satisfying (2.10) on $\Sigma\times(0,\tau_{1}),$ assume that
$|h|_{g}(x,\tau)$ $\leq\delta\,\,\text{on}\,\,\left\{(x,t)\in\Sigma\times[0,\tau_{1}]:\,\gamma e^{\tau+\tau_{0}}<\frac{|F|^{2}}{4}(x)\right\},$
$\sum_{l=0}^{3}|\nabla^{l}h|_{g}(x,0)$ $\leq\delta\,\,\text{on}\,\,x\in\left\{\gamma e^{\tau_{0}}\leq\frac{|F|^{2}}{4}\right\},\text{ and}$
$|\mathcal{Q}_{1}|_{g}(x,\tau))+|\nabla^{g}\mathcal{Q}_{1}|_{g}(x,\tau)$ $\leq\delta\,\,\text{on}\,\,\left\{(x,\tau)\in\Sigma\times[0,\tau_{1}]:\,\gamma e^{\tau+\tau_{0}}<\frac{|F|^{2}}{4}(x)\right\}.$
Then for $\delta<\delta_{0}(n,\Sigma,F)$ and $\tau_{0}>\tau(n,\Sigma,F,\gamma)$, there exists $C=C(n,\Sigma,F)<\infty$ such that
$\sum_{l=0}^{2}|\nabla^{g}h|_{g}(x,\tau)\leq C\delta\,\,\text{on}\,\,\left\{(x,\tau)\in\Sigma\times[0,\tau_{1}]:\,4\gamma e^{\tau+\tau_{0}}<\frac{|F|^{2}}{4}(x)\right\}.$
5. Barrier functions
In this section, we will use a barrier function argument to get the corresponding $C_{0}$ estimate. Since the original evolution equation (2.10) is highly nonlinear, to construct barriers, we have to find the correct operator first. We will consider similar barrier functions as in the Ricci flow case [Sto] while we need to handle the nonlinearity of mean curvature flows. Recall that in normal coordinates, the second order term in $\mathcal{Q}_{2}$ is
(5.1) $C^{ab}(F,\nabla^{g}F,h,\nabla^{g}h)\partial_{a}\partial_{b}h\cdot\textbf{n}_{F}:=\frac{\tilde{h}^{ab}\left(1-|\nabla h|^{2}\right)-\left(|\nabla h|^{2}+\tilde{h}^{cd}\partial_{c}h\,\partial_{d}h\right)g^{ab}}{1+\langle\textbf{n}_{F},\widetilde{\textbf{n}}\rangle}\partial_{a}\partial_{b}h\cdot\textbf{n}_{F}.$
We look at the lower order terms of $\mathcal{Q}_{2}$ in (2.10). Using (2.9), if we denote $C=C(F,\nabla F,\nabla^{2}F,h,\nabla h)$ to be any tensor depending on $F,\nabla F,\nabla^{2}F,h,\nabla h$, we get
$C_{3}h+C_{4}*\nabla h$
$=$ $C*(\nabla h\textbf{n}_{F}+hA)*(\nabla h\textbf{n}_{F}+hA)+C*\nabla h*\nabla h*A+C*hA*A$
$+hC*(\nabla h\textbf{n}_{F}+hA)*(\nabla h\textbf{n}_{F}+hA)$
$+(h+C)(C*\nabla h+C*\nabla h\textbf{n}_{F}+C*hA)*(\nabla h+C*\nabla h\textbf{n}_{F}+C*hA)$
$+(ch*A+C*\nabla h*\nabla h)(C*\nabla h*A+hC*\nabla A+hC*A*A)$
$+\nabla h*(h*\nabla A+A*\nabla h+h\nabla h*A*A+h^{2}A*\nabla A)*(C+h*A+C*\nabla h*\nabla h+C*hA),$
and we can simplify the expression to get
$C_{3}h+C_{4}*\nabla h=C*\nabla h*\nabla h+C*hA*A+C*\nabla h*hA+Ch*\nabla A.$
Thus, we know
$|C_{3}h+C_{4}*\nabla h|\leq C|\nabla h|^{2}+C|h||A|^{2}+C|h||\nabla A|,$
so we have
$\partial_{\tau}|h|^{2}$ $=2\langle h\,\textbf{n}_{F},\partial_{t}h\,\textbf{n}_{F}\rangle=2h(Lh+\mathcal{Q}_{1}+\mathcal{Q}_{2})$
$\leq\left(\Delta-\nabla_{\nabla^{g}\frac{|F|^{2}}{4}}\right)|h|^{2}-2|\nabla h|^{2}+\left(2|A|^{2}+1\right)h^{2}+C^{ab}\nabla_{a}\nabla_{b}h^{2}-2C^{ab}\nabla_{a}h\nabla_{b}h$
$\quad+2h\mathcal{Q}_{1}+h\left(C|\nabla h|^{2}+C|h||A|^{2}+C|h||\nabla A|\right).$
Using (5.1), given $|h|_{C_{1}}\leq\varepsilon<\varepsilon_{n},$ we have
$|C^{ab}|\leq C_{n}\left(|hA|_{C^{0}}+|\nabla h|^{2}_{C^{0}}\right),$
and
$\partial_{\tau}|h|^{2}$ $\leq\left(g^{ab}+C^{ab}\right)\nabla_{a}\nabla_{b}h^{2}-\nabla_{\nabla^{g}\frac{|F|^{2}}{4}}|h|^{2}+\left(C_{n}|A|^{2}+C_{n}|\nabla A|+1\right)h^{2}+2|h||\mathcal{Q}_{1}|.$
These imply that we can look at the operator
(5.2) $Qf:=\left(g^{ab}+C^{ab}\right)\nabla_{a}\nabla_{b}f-\nabla_{\nabla^{g}\frac{|F|^{2}}{4}}f+\left(C_{n}|A|^{2}+C_{n}|\nabla A|+1\right)f+2\sqrt{|f|}|\mathcal{Q}_{1}|.$
We now take $f=\frac{|F|^{2}}{4}.$ From the shrinker equation, we have $\Delta\frac{|F|^{2}}{4}-\nabla_{\nabla^{g}\frac{|F|^{2}}{4}}\frac{|F|^{2}}{4}=\frac{n}{2}-f$. Hence, we obtain
$\left(g^{ab}+C^{ab}\right)\nabla_{a}\nabla_{b}f^{k}-\nabla_{\nabla^{g}\frac{|F|^{2}}{4}}f^{k}=kf^{k-1}\left(\frac{n}{2}-f\right)+k(k-1)f^{k-1}\frac{|\nabla f|^{2}}{f}+C^{ab}\nabla_{a}\nabla_{b}f^{k}.$
As
(5.3) $\left|C^{ab}\nabla_{a}\nabla_{b}f^{k}\right|$ $=\left|kf^{k-1}C^{ab}\nabla_{a}\nabla_{b}f+k(k-1)C^{ab}f^{k-2}\nabla_{a}f\nabla_{b}f\right|$
$\leq\left|kf^{k-1}\right|\cdot\left|C^{ab}\right|\left(|A_{ab}|f^{\frac{1}{2}}+(k-1)|F_{a}||F_{b}|\right),$
similarly as in [Sto], we take
$u=\begin{cases}e^{(-k+1)(\tau+\tau_{0})}\left(Df^{k}-Bf^{k-1}\right)&\text{on }\left\{(x,\tau)\in\Sigma\times[0,\infty):\Gamma<f(x)<\gamma_{1}e^{\tau+\tau_{0}}\right\}\\ e^{\tau+\tau_{0}}\left(a-Bf^{-1}\right)&\text{on }\left\{(x,\tau)\in\Sigma\times[0,\infty):\gamma_{2}e^{\tau+\tau_{0}}<f(x)<\Gamma e^{\tau+\tau_{0}}\right\}\end{cases}.$
In general, using (5.3) and $0\leq\frac{|\nabla f|^{2}}{f}\leq 1,$ we calculate
$e^{(m-1)(\tau+\tau_{0})}(\partial_{\tau}-Q)u$
$=$ $e^{(m-1)(\tau+\tau_{0})}(\partial_{\tau}-Q)\left(e^{(-m+1)(\tau+\tau_{0})}\left(Df^{m}-Bf^{m-1}\right)\right)$
$=$ $-\frac{n}{2}\left(mDf^{m-1}-(m-1)Bf^{m-2}\right)+Bf^{m-1}-(m(m-1)Df^{m-1}-(m-1)(m-2)Bf^{m-2})\frac{|\nabla f|^{2}}{f}$
$-C^{ab}\nabla_{a}\nabla_{b}\left(Df^{m}-Bf^{m-1}\right)-\left(C_{n}|A|^{2}+C_{n}|\nabla A|\right)\left(Df^{m}-Bf^{m-1}\right)-e^{(m-1)(\tau+\tau_{0})}2\sqrt{|u|}|\mathcal{Q}_{1}|$
$\geq$ $f^{m-1}\left(B+\frac{n(m-1)B}{2f}-\frac{nmD}{2}-m(m-1)D-Dm\left|C^{ab}\right|\left|A_{ab}\right|f^{\frac{1}{2}}-Dm(m-1)\left|C^{ab}\right|\right.$
$\left.-B(m-1)\left|C^{ab}\right|\left|A_{ab}\right|f^{-\frac{1}{2}}-\frac{B(m-1)(m-2)\left|C^{ab}\right|}{f}-\left(C_{n}|A|^{2}+C_{n}|\nabla A|\right)(Df+B))\right)$
$-e^{(m-1)(\tau+\tau_{0})}2\sqrt{|e^{(-m+1)(\tau+\tau_{0})}(Df^{m}-Bf^{m-1})|}|\mathcal{Q}_{1}|.$
Therefore, as $\mathcal{Q}_{1}$ is 0 on $\left\{(x,\tau)\in\Sigma\times[0,\infty):\Gamma e^{\tau_{0}}<f(x)<\gamma_{1}e^{\tau+\tau_{0}}\right\}$, there exist constants $\varepsilon_{n},\Gamma,$ and $\gamma_{1}$ such that given
$|h|_{C_{1}}\leq\varepsilon<\varepsilon_{n},\quad B-D\left(m(m+\frac{n}{2})+C_{n}\sup_{\Sigma}\left(f|A|^{2}+f|\nabla A|\right)\right)\geq w>0$
$\Gamma\geq\Gamma(n,B,m,w),\text{ and }\gamma_{1}\leq\gamma_{1}(n,\Sigma,F,B,D,m,w),$
$u$ satisfies $(\partial_{t}-Q)u\geq 0$ on $\left\{(x,\tau)\in\Sigma\times[0,\infty):\Gamma e^{\tau_{0}}<f(x)<\gamma_{1}e^{\tau+\tau_{0}}\right\}$.
Moreover, if
$\mathcal{Q}_{1}\leq\frac{C_{0}}{\Gamma}e^{-\frac{\tau+\tau_{0}}{2}},$
then on $\left\{(x,\tau)\in\Sigma\times[0,\infty):\gamma_{2}e^{(\tau+\tau_{0})}<f(x)<\Gamma e^{(\tau+\tau_{0})}\right\},$
$e^{-(\tau+\tau_{0})}2\sqrt{|e^{(\tau+\tau_{0})}(a-Bf^{-1})|}\left|\mathcal{Q}_{1}\right|\leq\frac{2\sqrt{a}}{e^{(\tau+\tau_{0})}\Gamma}\leq 2\frac{\sqrt{a}}{f}.$
Plugging in the expression, there exists $\tau_{0}$ such that given
$|h|_{C_{1}}\leq\varepsilon<\varepsilon_{n},\quad B-a\left(C_{n}\sup_{\Sigma}\left(f|A|^{2}+f|\nabla A|\right)\right)-C_{0}\sqrt{a}\geq w>0,\text{ and}$
$\tau_{0}\geq\tau_{0}(n,\Sigma,F,C_{0},B,\gamma_{2},\varepsilon_{n}),$
$u$ satisfies $(\partial_{\tau}-Q)u\geq 0$ on $\left\{(x,\tau)\in\Sigma\times[0,\infty):\gamma_{2}e^{(\tau+\tau_{0})}<f(x)<\Gamma e^{(\tau+\tau_{0})}\right\}$.
Therefore, we have the following estimates.
[…]
Given $|h|_{C_{1}}\leq\varepsilon<\varepsilon_{n},$ we have $(\partial_{\tau}-Q)h^{2}\leq 0,$ where $Q$ is given by (5.2). Moreover, if $\mathcal{Q}_{1}$ satisfies
$\left|\mathcal{Q}_{1}\right|\begin{cases}=0&\text{on }\left\{(x,\tau)\in\Sigma\times[0,\infty):\Gamma e^{\tau_{0}}<f(x)<\gamma_{1}e^{(\tau+\tau_{0})}\right\}\\ \leq\frac{C_{0}}{\Gamma}e^{-\frac{(\tau+\tau_{0})}{2}}&\text{on }\left\{(x,\tau)\in\Sigma\times[0,\infty):\gamma_{2}e^{(\tau+\tau_{0})}<f(x)<\Gamma e^{(\tau+\tau_{0})}\right\}\end{cases},$
then
$u:=\begin{cases}e^{(-k+1)(\tau+\tau_{0})}(Df^{k}-Bf^{k-1})&\text{on }\left\{(x,\tau)\in\Sigma\times[0,\infty):\Gamma e^{\tau_{0}}<f(x)<\gamma_{1}e^{(\tau+\tau_{0})}\right\}\\ e^{(\tau+\tau_{0})}(a-Ef^{-1})&\text{on }\left\{(x,\tau)\in\Sigma\times[0,\infty):\gamma_{2}e^{(\tau+\tau_{0})}<f(x)<\Gamma e^{(\tau+\tau_{0})}\right\}\end{cases}$
satisfies $(\partial_{t}-Q)u\geq 0$ on its domain if
$|h|_{C_{1}}\leq\varepsilon<\varepsilon_{n},\quad B-D\left(m(m+\frac{n}{2})+C_{n}\sup_{\Sigma}\left(f|A|^{2}+f|\nabla A|\right)\right)\geq w>0$
$\Gamma\geq\Gamma(n,B,m,w),\quad\gamma_{1}\leq\gamma_{1}(n,\Sigma,F,B,D,m,w),$
$E-a\left(C_{n}\sup_{\Sigma}\left(f|A|^{2}+f|\nabla A|\right)\right)-C_{0}\sqrt{a}\geq w>0,\text{ and }\tau_{0}\geq\tau_{0}(n,\Sigma,F,C_{0},B,\gamma_{2},\varepsilon_{n}).$
This proposition seems very technical, but it enables us to easily control the $C^{0}$ norm of h using this barrier function. By combining Proposition 4, Lemma 4, Lemma 4, and Lemma 4, we can prove the global $C^{2}$-improvement theorem following the same argument in [Sto]. Note that here we need $C^{1}$-bounds instead of $C^{0}$-bounds in the mean curvature flow setting.
[…]
Given $\Gamma>0$ and $h$ satisfying (2.10) on $\Sigma\times(0,\tau_{1})$ and $\tau=-\log(1-t)$, assume
$\|h\|_{C^{1}(\Sigma\times[0,\tau_{1}])}$ $<\varepsilon,$
$h(0)$ $=\eta\left(\frac{1}{\gamma_{0}e^{\tau_{0}}}f(x)\right)\sum_{i=1}^{m}p_{i}f_{i}\text{ for }|\mathbf{p}|\leq\bar{p}e^{\lambda_{*}\tau_{0}}$
$\|h(\tau)\|_{L_{W}^{2}(\Sigma)}$ $\leq\mu e^{\lambda_{*}(\tau+\tau_{0})}\text{ for all }\tau\in[0,\tau_{1}],$
$\text{supp}(\mathcal{Q}_{1}(\cdot,\tau))$$\subseteq\left\{\frac{\Gamma_{0}e^{\frac{(\tau+\tau_{0})}{2}}}{2}\leq f\leq\Gamma_{0}e^{\frac{(\tau+\tau_{0})}{2}}\right\},$
$|\mathcal{Q}_{1}(\cdot,\tau)|$ $\leq C\left(\frac{1}{\Gamma_{0}e^{(\tau+\tau_{0})}}\right)^{\frac{1}{2}},\text{ and}$
$|\nabla\mathcal{Q}_{1}(\cdot,\tau)|$ $\leq C\left(\frac{1}{\Gamma_{0}e^{\tau+\tau_{0}}}\right).$
Then for any $\delta<\varepsilon$, $\Gamma>\Gamma_{0}(n,\Sigma,F)$, $\varepsilon<\varepsilon_{0}(n,\Sigma,F,\Gamma)$, $|\mathbf{p}|<C(n,\Sigma,F,\lambda_{*})$, $\mu<\mu_{0}(n,\Sigma,F,\lambda_{*})$, $\gamma<\gamma_{0}(n,\Sigma,F,\lambda_{*},\delta),$ and $\tau_{0}>C(n,\Sigma,F,\lambda_{*},\Gamma,\bar{p},\varepsilon,\delta)$, there exist $W=W(n,\Sigma,F,\lambda_{*})<\infty$ and $W^{\prime}=W^{\prime}(n,\Sigma,F)<\infty$ such that
$|h|+|\nabla h|+|\nabla^{2}h|$ $\leq W\left(\frac{|F|^{2}}{4}\right)^{|\lambda_{*}|}e^{\lambda_{*}{(\tau+\tau_{0})}}\text{ for }(x,\tau)\in\Sigma\times[0,\tau_{1}],\text{ and}$
$|h|+|\nabla h|+|\nabla^{2}h|$ $\leq W^{\prime}\delta\text{ for }(x,\tau)\in\Sigma\times[0,\tau_{1}].$
6. Ważewski’s box argument
We are going to finish the proof of the main theorem in this section. For the concerned asymptotically conical shrinker $\Sigma,$ take $\tau_{0},\Gamma_{0}>0$ sufficiently large. From Section 3, we can construct $M=M_{\Sigma,\tau_{0},\Gamma_{0}},$ a doubling of a large compact set $\{f(x)\leq 2e^{\tau_{0}}\Gamma_{0}\}$ of $\Sigma$. Based on Proposition 2.3, we take an $L^{2}_{W}$-orthonormal eigenbasis $\{h_{i}\}_{i\in\mathbb{N}}$ of $L$ on $\Sigma$ and the corresponding eigenvalues $\{\lambda_{i}\}_{i\in\mathbb{N}}$. Take $\lambda_{*}<0$ and $m$ such that $\lambda_{m+1}<\lambda_{*}<\lambda_{m}$.
We fix a compactly supported bump function $\eta:(0,\infty)\to[0,1]$ such that $\eta(x)=1$ for $x<\frac{1}{2}$ and $\eta(x)=0$ for $x>1$. Take any $\mathbf{p}=(p_{1},p_{2},\cdots,p_{m})\in\mathbb{R}^{m},$ and then we can perturb the immersion of $M$ into $\mathbb{R}^{n+1}$ such that
$F_{\mathbf{p}}(x,0)=F_{M}(x)+\eta\left(\frac{1}{\gamma_{0}e^{\tau_{0}}}f(x)\right)\sum_{i=1}^{m}p_{i}f_{i}\cdot\textbf{n}_{F}(x).$
We start the mean curvature flow from $F_{\mathbf{p}}(x,0)$. We denote the family of immersions of mean curvature flows by $\{F(x,t)\}_{t\in[0,T(\mathbf{p})]}$ and transplant it back to the original shrinker as
$\hat{F}_{\mathbf{p}}(x,t)=\eta\left(\frac{1}{\Gamma_{0}e^{\tau_{0}}}f(x)\right)F_{\mathbf{p}}(x,t)+\left(1-\eta\left(\frac{1}{\Gamma_{0}e^{\tau_{0}}}f(x)\right)\right)F(x,t),$
where $F(x,t)$ corresponds to the family of immersions of mean curvature flows of the shrinker $\Sigma$. From the discussion in Section 2.2 we can take $F_{0}=\hat{F}_{\mathbf{p}}$. Then we get
$\frac{1}{\sqrt{1-t}}\hat{F}_{\mathbf{p}}(\varphi_{t}(x),t)=F(x)+h_{\mathbf{p}}(x,t)\textbf{n}_{F}(x)$
and
$d\hat{F}_{\mathbf{p}}(\partial_{t}\varphi_{t}(x))=\frac{\left\langle(\partial_{t}\hat{F}_{\mathbf{p}})(\varphi_{t}(x),t)+\frac{1}{2\sqrt{1-t}}F(x),\textbf{n}_{\hat{F}_{\mathbf{p}}}(\varphi_{t}(x))\right\rangle}{\left\langle\textbf{n}_{F}(x),\textbf{n}_{\hat{F}_{\mathbf{p}}}(\varphi_{t}(x))\right\rangle}\textbf{n}_{F}(x)-\left((\partial_{t}\hat{F}_{\mathbf{p}})(\varphi_{t}(x),t)+\frac{1}{2\sqrt{1-t}}F(x)\right).$
In particular, we have
(6.1) $\left|d\hat{F}_{\mathbf{p}}(\partial_{t}\varphi_{t}(x))\right|\leq\frac{\left|(\partial_{t}\hat{F}_{\mathbf{p}})(\varphi_{t}(x),t)+\frac{1}{2\sqrt{1-t}}F(x)\right|}{\left|\left\langle\textbf{n}_{F}(x),\textbf{n}_{\hat{F}_{\mathbf{p}}}(\varphi_{t}(x))\right\rangle\right|}.$
From (2.10), the function $h=h_{\mathbf{p}}$ satisfies the equation
$\frac{\partial h}{\partial\tau}=$ $Lh+\frac{{\sqrt{1-t}}\langle(\partial_{t}\hat{F}_{\mathbf{p}}-\Delta^{g_{0}}\hat{F}_{\mathbf{p}})(\varphi_{t}(x),t),\textbf{n}_{F_{1}}(x)\rangle}{\langle\textbf{n}_{F}(x),\textbf{n}_{F_{1}}(x)\rangle}$
$+C_{3}(F,\nabla F,h,\nabla h)*h+C_{4}(F,\nabla F,h,\nabla h)*\nabla h+C_{5}(F,\nabla F,h,\nabla h)*\nabla^{2}h$
$=$ $Lh+\mathcal{Q}_{1}+\mathcal{Q}_{2},$
for $\tau=-\log(1-t),$ where
$\mathcal{Q}_{1}=$ ${\sqrt{1-t}}\langle(\partial_{t}\hat{F}_{\mathbf{p}}-\Delta^{g_{0}}\hat{F}_{\mathbf{p}})(\varphi_{t}(x),t),\textbf{n}_{F_{1}}(x)\rangle,\text{ and}$
$\mathcal{Q}_{2}=$ $C_{3}(F,\nabla F,h,\nabla h)*h+C_{4}(F,\nabla F,h,\nabla h)*\nabla h+C_{5}(F,\nabla F,h,\nabla h)*\nabla^{2}h$
$+{\sqrt{1-t}}\left\langle(\partial_{t}\hat{F}_{\mathbf{p}}-\Delta^{g_{0}}\hat{F}_{\mathbf{p}})(\varphi_{t}(x),t),\textbf{n}_{F_{1}}(x)\right\rangle\left(\frac{1}{{\langle\textbf{n}_{F}(x),\textbf{n}_{F_{1}}(x)\rangle}}-1\right).$
In normal coordinates, the second order term in $\mathcal{Q}_{2}$ corresponds to
(6.2) $\frac{\tilde{h}^{ab}\left(1-|\nabla h|^{2}\right)-\left(|\nabla h|^{2}+\tilde{h}^{cd}\partial_{c}h\,\partial_{d}h\right)g^{ab}}{\left(1+\left\langle\textbf{n}_{F},\widetilde{\textbf{n}}\right\rangle\right)^{2}}\partial_{a}\partial_{b}h\cdot\textbf{n}_{F}.$
Given any $h\in L^{2}_{W}(\Sigma)$, we denote the stable and unstable projections by
$h_{u}$ $:=\sum_{j=1}^{m}\langle h,h_{j}\rangle_{L^{2}_{W}}h_{j},\text{ and}$
$h_{s}$ $:=\sum_{j=m+1}^{\infty}\langle h,h_{j}\rangle_{L^{2}_{W}}h_{j}=h-h_{u}.$
Similar to the Ricci flow case [Sto], we define the box
$\mathcal{B}[\lambda_{*},\mu_{u},\mu_{s},\varepsilon_{0},\varepsilon_{1},\varepsilon_{2},t_{1}]$
to be the set which consists of all the functions $h\colon\Sigma\times[0,-\log(1-t_{1})]\to\mathbb{R}$ satisfying the following conditions.
(1) $h(\cdot,\tau)$ is compactly supported in $\Sigma$ for any $\tau\in[0,-\log(1-t_{1})].$
(2) For $\tau\in[0,-\log(1-t_{1})],$
$\|h_{u}(\cdot,\tau)\|_{L^{2}_{W}}$ $\leq\mu_{u}e^{\lambda_{*}(\tau+\tau_{0})},\text{ and}$
$\|h_{s}(\cdot,\tau)\|_{L^{2}_{W}}$ $\leq\mu_{s}e^{\lambda_{*}{(\tau+\tau_{0})}}.$
(3) For $\tau\in[0,-\log(1-t_{1})],$
$|h|_{\bar{g}}$ $\leq\varepsilon_{0},$
$|\nabla h|_{\bar{g}}$ $\leq\varepsilon_{1},\text{ and}$
$|\nabla^{2}h|_{\bar{g}}$ $\leq\varepsilon_{2}.$
Our main goal here is to prove that for $\gamma_{0},|\mathbf{p}|<<1$ and $\Gamma_{0},\tau_{0}>>1$, $h_{\mathbf{p}}$ can only leave $\mathcal{B}$ from the unstable side.
[…]
Given $\gamma_{0},|\mathbf{p}|\leq\bar{p}e^{\lambda_{*}\tau_{0}}<<1$ and $\Gamma_{0},\tau_{0}>>1$, for $h_{\mathbf{p}}$ constructed above, assume $h_{\mathbf{p}}\notin\mathcal{B}[\lambda_{*},\mu_{u},\mu_{s},\varepsilon_{0},\varepsilon_{1},\varepsilon_{2},1].$ If we denote the largest existence time for $F_{\mathbf{p}}(x,\cdot)$ by $T(\mathbf{p})$ and take the largest time $t(\mathbf{p})$ such that $h_{\mathbf{p}}\in\mathcal{B}=\mathcal{B}[\lambda_{*},\mu_{u},\mu_{s},\varepsilon_{0},\varepsilon_{1},\varepsilon_{2},t(\mathbf{p})]$, then we have
$\|h_{\mathbf{p},u}(\cdot,-\log(1-t(\mathbf{p})))\|_{L^{2}_{W}}=\mu_{u}e^{\lambda_{*}(\tau_{0}-\log(1-t(\mathbf{p})))}.$
Proof.
From the standard ordinary differential equation and parabolic theories, we have that for $|\mathbf{p}|<<1$, we can get $t(\mathbf{p})>0$. From the interior estimate and the pseudolocality in [EH, INS], on $M\setminus\left\{f\leq\frac{\Gamma_{0}e^{\tau_{0}}}{2}\right\}$, $F_{\mathbf{p}}(x,t)$ satisfies $|\nabla^{m}A|\leq C_{m}(\frac{1}{\Gamma_{0}e^{\tau_{0}}})^{\frac{m+1}{2}}$ for $t\in[0,2]$. Thus, if $h_{\mathbf{p}}\in\mathcal{B}$, we have
$|(\partial_{t}\hat{F}_{\mathbf{p}}-\Delta^{g_{0}}\hat{F}_{\mathbf{p}})|(x,t)$ $\leq C\left(\frac{1}{\Gamma_{0}e^{\tau_{0}}}\right)^{\frac{1}{2}},\text{ and}$
$|\nabla(\partial_{t}\hat{F}_{\mathbf{p}}-\Delta^{g_{0}}\hat{F}_{\mathbf{p}})|(x,t)$ $\leq C\left(\frac{1}{\Gamma_{0}e^{\tau_{0}}}\right).$
Hence, if we denote $\tau=-\log(1-t)$,
$\text{supp}(\mathcal{Q}_{1}(\cdot,\tau))$ $=\text{supp}\left({\sqrt{1-t}}\left\langle(\partial_{t}\hat{F}_{\mathbf{p}}-\Delta^{g_{0}}\hat{F}_{\mathbf{p}})(\varphi_{t}(\cdot),t),\textbf{n}_{F_{1}}(\cdot)\right\rangle\right)\subseteq\left\{\frac{\Gamma_{0}e^{\frac{(\tau+\tau_{0})}{2}}}{2}\leq f\leq\Gamma_{0}e^{\frac{(\tau+\tau_{0})}{2}}\right\},$
$|\mathcal{Q}_{1}(\cdot,t)|$ $\leq C\left(\frac{1}{\Gamma_{0}e^{\tau+\tau_{0}}}\right)^{\frac{1}{2}},\text{ and }|\nabla\mathcal{Q}_{1}(\cdot,t)|\leq C\left(\frac{1}{\Gamma_{0}e^{\tau+\tau_{0}}}\right).$
From Theorem 5, given $\tau_{0}>>1,$ we can improve the $C^{2}$ bounds such that
$\|h\|_{C^{2}(\Sigma\times[0,\tau_{1}])}\leq W^{\prime}\delta<\frac{1}{2}\min\{\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}\},\text{ and}$
$\|h\|_{C^{2}(\Sigma\times[0,\tau_{1}])}\leq W\left(\frac{|F|^{2}}{4}\right)^{|\lambda_{*}|}e^{\lambda_{*}(\tau+\tau_{0})}.$
From the equation, given $|h|_{C^{2}}<\varepsilon_{0}(n)$,
$|\mathcal{Q}_{2}|\leq C\left(|h|+|\nabla h|+|\nabla^{2}h|\right)^{2}\leq C\min\left\{W\left(\frac{|F|^{2}}{4}\right)^{|\lambda_{*}|}e^{\lambda_{*}(\tau+\tau_{0})},\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}\right\}.$
Thus, given $\tau_{0}>>1$,
(6.3) $\|\mathcal{Q}_{1}\|_{L^{2}_{W}}+\|\mathcal{Q}_{2}\|_{L^{2}_{W}}\leq CW^{2}e^{2\lambda_{*}(\tau+\tau_{0})}\leq Ce^{2\lambda_{*}(\tau+\tau_{0})}.$
From the definition of $h_{\mathbf{p}}$, the polynomial decay of eigenfunctions (Theorem 2.4), and the exponential decay of the weight, for $|\mathbf{p}|\leq\bar{p}e^{\lambda_{*}\tau_{0}}$, we have
(6.4) $\left|\langle h_{\mathbf{p}}(\cdot,0),h_{j}\rangle_{L^{2}_{W}}-p_{j}\right|$ $\leq Ce^{-\frac{\gamma_{0}}{100}e^{\tau_{0}}}\text{ for }j\leq m,\text{ and}$
$\|h_{\mathbf{p},s}\|_{L^{2}_{W}}$ $\leq Ce^{-\frac{\gamma_{0}}{100}e^{\tau_{0}}}.$
From (6.3) and the equation $\partial_{\tau}h=Lh+\mathcal{Q}_{1}+\mathcal{Q}_{2}$,
(6.5) $\frac{d}{d\tau}\|e^{-\lambda_{*}(\tau+\tau_{0})}h_{\mathbf{p},u}\|_{L^{2}_{W}}$ $\geq(\lambda_{m}-\lambda_{*})\|e^{-\lambda_{*}(\tau+\tau_{0})}h_{\mathbf{p},u}\|_{L^{2}_{W}}-Ce^{\lambda_{*}\tau_{0}},\text{ and}$
$\frac{d}{d\tau}\|e^{-\lambda_{*}(\tau+\tau_{0})}h_{\mathbf{p},s}\|_{L^{2}_{W}}$ $\leq(\lambda_{m+1}-\lambda_{*})\|e^{-\lambda_{*}(\tau+\tau_{0})}h_{\mathbf{p},u}\|_{L^{2}_{W}}+Ce^{\lambda_{*}\tau_{0}}.$
In particular, given $\tau_{0}>C(n,\mu_{s},\mu_{u}),$
$\|h_{\mathbf{p},s}(\cdot,-\log(1-t(\mathbf{p})))\|_{L^{2}_{W}}<\mu_{s}e^{\lambda_{*}(\tau_{0}-\log(1-t(\mathbf{p})))},$
and thus, from the definition of the box, ww obtain
$\|h_{\mathbf{p},u}(\cdot,-\log(1-t(\mathbf{p})))\|_{L^{2}_{W}}=\mu_{u}e^{\lambda_{*}(\tau_{0}-\log(1-t(\mathbf{p})))}.$
Moreover, from (6.5), for $\tau>-\log(1-t(\mathbf{p}))$,
$\|h_{\mathbf{p},u}(\cdot,\tau)\|_{L^{2}_{W}}>\mu_{u}e^{\lambda_{*}(\tau+\tau_{0})}.$
This completes the proof. ∎
Next, we prove that we can choose a suitable perturbation such that the flow remains in the box until time $t=1$. This is based on a topological argument.
[…]
Given $\gamma_{0},$ $\bar{p}e^{\lambda_{*}\tau_{0}}<<1$ and $\Gamma_{0},\tau_{0}>>1$, there exists $\mathbf{p}\in\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}$ such that
(6.6) $h_{\mathbf{p}}\in\mathcal{B}[\lambda_{*},\mu_{u},\mu_{s},\varepsilon_{0},\varepsilon_{1},\varepsilon_{2},1].$
Proof.
As in the Ricci flow case, we define $\mathcal{F}:\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}\to\mathbb{R}^{m}$ by
$\mathcal{F}(\mathbf{p})=\left(\left\langle h_{\mathbf{p}}(t(\mathbf{p})),h_{i}\right\rangle_{L^{2}_{W}}\right)_{i=1,2,\cdots,m}.$
If we take $\mu_{u}<\frac{\bar{p}}{2}$, from (6.4), the map
$\mathcal{F}|_{\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}\setminus\overline{B_{2\mu_{u}e^{\lambda_{*}\tau_{0}}}}}\colon\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}\setminus\overline{B_{2\mu_{u}e^{\lambda_{*}\tau_{0}}}}\to\mathbb{R}^{m}\setminus\{0\}$
is homotopic to the identity through $\mathcal{F}_{s}:=s\mathcal{F}+(1-s){\rm Id}\colon\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}\setminus\overline{B_{2\mu_{u}e^{\lambda_{*}\tau_{0}}}}\to\mathbb{R}^{m}\setminus\{0\}$.
On the other hand, if
$h_{\mathbf{p}}\notin\mathcal{B}[\lambda_{*},\mu_{u},\mu_{s},\varepsilon_{0},\varepsilon_{1},\varepsilon_{2},1]$
for any $\mathbf{p}\in\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}},$ then Proposition 6 implies that $|\mathcal{F}(\mathbf{p})|>0$ for $\mathbf{p}\in\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}$. This shows that, in particular,
$\mathcal{F}|_{\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}\setminus\overline{B_{2\mu_{u}e^{\lambda_{*}\tau_{0}}}}}:\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}\setminus\overline{B_{2\mu_{u}e^{\lambda_{*}\tau_{0}}}}\to\mathbb{R}^{m}\setminus\{0\}$
is null-homotopic.
Hence, we derive a contraction. This proves that there exists $\mathbf{p}\in\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}$ such that
(6.7) $h_{\mathbf{p}}\in\mathcal{B}[\lambda_{*},\mu_{u},\mu_{s},\varepsilon_{0},\varepsilon_{1},\varepsilon_{2},1],$
completing the proof. ∎
We can prove our main result (Theorem 1).
[…]
Given an asymptotically conical shrinker $\Sigma^{n}\subseteq\mathbb{R}^{n+1}$, there exists a closed hypersurface $M\subseteq\mathbb{R}^{n+1}$ such that the mean curvature flow $M_{t}$ starting from $M$ develops a type I singularity at time $1$ at the origin modeled on $\Sigma$.
Proof.
For any $\mathbf{p},$ if we blow up the mean curvature flow starting from the corresponding perturbation at $({\bf 0},1)$, the family of rescaled hypersurfaces is the graph of $h_{\mathbf{p}}$ over $\Sigma$. From Proposition 6, there exists a vector $\mathbf{p}$ such that $h_{\mathbf{p}}\in\mathcal{B}[\lambda_{*},\mu_{u},\mu_{s},\varepsilon_{0},\varepsilon_{1},\varepsilon_{2},1].$ Based on Theorem 5 and Proposition 4, $h_{\mathbf{p}}$ converges to 00 in $C^{\infty}_{loc}(\Sigma)$. The global $C^{2}$ bound of $h_{\mathbf{p}}$ guarantees that the singularity is of type I. This finishes the proof. ∎

### v2

@@
 Abstract.
-In this paper, we prove that for any asymptotically conical self-shrinker, there exists a closed hypersurface such that the mean curvature flow starting from it develops a singularity modeled on the given shrinker.
-The main technique is the Ważewski box argument, used by Stolarski [Sto] in the proof of the corresponding theorem in the Ricci flow case.
+In this paper, we prove that for any asymptotically conical self-shrinker, there exists an embedded closed hypersurface such that the mean curvature flow starting from it develops a singularity modeled on the given shrinker.
+The main technique is the Ważewski box argument, used by Stolarski in the proof of the corresponding theorem in the Ricci flow case.
+As a corollary, our construction, combined with the works of Angenent–Ilmanen–Velázquez and Chodosh–Daniels-Holgate–Schulze, implies the existence of fattening level set flows starting from smooth embedded closed hypersurfaces.
+This addresses a question posed by Evans–Spruck and De Giorgi.
 1. Introduction
@@
 More precisely, given a mean curvature flow $M_{t}$ developing a spacetime singularity at $(\textbf{0},0)\in\mathbb{R}^{n+k}\times\mathbb{R},$ we can consider a sequence of rescaled mean curvature flows $M^{i}_{t}:=\lambda_{i}M_{\lambda_{i}^{-2}t}$ using any sequence $\lambda_{i}\to\infty.$ Based on Huisken’s monotonicity [H90] and Brakke’s compactness [B78], White [W94] and Ilmanen [I95] proved that $M_{t}^{i}$ weakly converges to a self-similar limit flow, called a tangent flow, after replacing $\lambda_{i}$’s with a subsequence.
-When the limiting flow is smooth, it is generated by a self-shrinker, a submanifold satisfying $\textbf{H}=-\frac{x^{\perp}}{2}.$ Uniqueness of tangent flows, a fundamental issue for singularity analysis, has been discussed for certain shrinkers [S14, CM15, CM19, Z20, CS, LSS, LZ, CSSZ].
+When the limit flow is smooth, it is generated by a self-shrinker, a submanifold satisfying $\textbf{H}=-\frac{x^{\perp}}{2}.$ Uniqueness of tangent flows, a fundamental issue for singularity analysis, has been discussed for certain shrinkers [S14, CM15, CM19, Z20, CS, LSS, S23, LZ, CSSZ].
 For related references on uniqueness of tangent cones of minimal submanifolds, one can see [AA, S83, H97, S08, CM14].
@@
 […]
-Given an asymptotically conical shrinker $\Sigma^{n}\subset\mathbb{R}^{n+1}$, there exists a closed embedded hypersurface $M\subset\mathbb{R}^{n+1}$ such that the mean curvature flow $M_{t}$ starting from $M$ develops a type I singularity at time $1$ at the origin modeled on $\Sigma$.
+Given an asymptotically conical shrinker $\Sigma^{n}\subset\mathbb{R}^{n+1}$, there exists an embedded closed hypersurface $M\subset\mathbb{R}^{n+1}$ such that the mean curvature flow $M_{t}$ starting from $M$ develops a type I singularity at time $1$ at the origin modeled on $\Sigma$.
 This confirms the belief in the existence of closed mean curvature flows having asymptotically conical tangent flows.
-It then has some applications based on earlier works.
-For example, Sun-Xue [SX2] proved that given a closed mean curvature flow with an asymptotically conical singularity, after a generic initial perturbation, the perturbed mean curvature flow avoids the asymptotically conical singularity.
-Their theorem suggests that it is hard to see an asymptotically conical singularity in closed mean curvature flows.
-In fact, it is expected and has been proven in the surface case (cf.
-[CCMS, CCS, SX1, SX2, BK23]) that all non-cylindrical singularities are not generic.
+This finding has applications, drawing on previous research.
+For instance, Chodosh–Choi–Mantoulidis–Schulze [CCMS] proved that for a closed mean curvature flow with an asymptotically conical singularity, a generic initial perturbation ensures the perturbed flow avoids the asymptotically conical singularity.
+Their theorem indicates that encountering an asymptotically conical singularity in closed mean curvature flows is unlikely.
+Indeed, it is anticipated that all non-cylindrical singularities are not generic, a conjecture confirmed for surfaces [CCMS, CCS, BK23].
+Another method addressing this issue, based on local dynamics, was explored by Sun–Xue [SX1, SX2].
+Another immediate consequence of Theorem 1 is noteworthy.
+In [ES], Evans–Spruck conjectured that any level set flow starting from a smooth hypersurface does not fatten (cf.
+the final sentence in [ES]).
+Based on the construction by Angenent–Ilmanen–Velázquez [AIV] and the fattening criterion proven by Chodosh–Daniels-Holgate–Schulze [CDS], Theorem 1 implies the following corollary.
+It provides a negative answer to the question raised by Evans–Spruck (cf.
+[De, I94, AIC]) by showing that fattening does not necessarily happen instantaneously.
+Corollary 1.1.
+There exists a fattening level set flow starting from an $n$-dimensional smooth embedded closed hypersurface for $n\in\left\{3,4,5,6\right\}.$ In particular, the flow remains smooth for a short time before fattening occurs.
+In [W02], White announced a result joint with Ilmanen about the existence of a compact smooth embedded surface in $\mathbb{R}^{3}$ for which the uniqueness of enhanced varifold solutions of mean curvature evolution fails.
+He described the proof ideas in [W02, Section 4], which are completely different from our approach to Corollary 1.1.
+Combining their result with Corollary 1.1, we now have $n$-dimensional embedded closed examples for $n\in\left\{2,3,4,5,6\right\}.$
 The proof of Theorem 1 follows the outline of the argument of the corresponding construction in the Ricci flow case by Stolarski [Sto].
@@
 In Sections 4 and 5, we talk about the interior estimates and constructions of barrier functions to get the preservation and the improvement of $C^{2}$ bounds.
-In Section 6, we use the Ważewski box argument to finish the proof.
+In Section 6, we use the Ważewski box argument to finish the proof of Theorem 1 and prove Corollary 1.1.
 Acknowledgement
@@
 Bill Minicozzi for their constant support and several inspiring conversations.
-They would also like to thank Max Stolarski for helpful conversations.
+They would like to thank Letian Chen, Otis Chodosh, Alec Payne, Max Stolarski, and Ao Sun for valuable discussions, and to Alec Payne for bringing the work of Angenent, Ilmanen, and Velázquez to their attention and suggesting the application regarding fattening.
 During the project, Lee was partially supported by NSF Grant DMS 2005345 and Zhao was supported by NSF Grant DMS 1812142 and NSF Grant DMS 1811267.
@@
 Let $\Sigma$ be an asymptotically conical shrinker in $\mathbb{R}^{n+1}.$ In this paper, this means that its asymptotic cone
-(2.1) $C=\lim\limits_{t\to 0}\sqrt{-t}\Sigma$
+(2.1) $\mathcal{C}=\lim\limits_{t\to 0}\sqrt{-t}\Sigma$
 is a regular cone, i.e., a cone whose singular set consists of the origin. This immediately implies a curvature bound
-(2.2) $\left|\nabla^{i}A\right|\leq\frac{C_{j}}{|x|^{j+1}}$
+(2.2) $\left|\nabla^{j}A\right|\leq\frac{C_{j}}{|x|^{j+1}}$
 for some $C_{j}=C_{j}(\Sigma)<\infty.$ One can see [W14] for a proof of this.
@@
 In [CS], it is also shown that $L\colon H^{2}_{W}\left(\Sigma\right)\to L^{2}_{W}(\Sigma)$ is a bounded operator.
-Hence, elliptic theories implies that the spectrum of $L$ behaves as nicely as we expected.
+Hence, elliptic theories imply that the spectrum of $L$ behaves as nicely as we expect.
 Proposition 2.3.
 There exists an $L^{2}_{W}$-orthonormal basis $\{h_{i}\}_{i\in\mathbb{N}}$ of $L^{2}_{W}(\Sigma)$ such that for all $i\in\mathbb{N},$ $h_{i}$ is an eigenfunction of $\mathcal{L}$ with eigenvalue $\lambda_{i},$ that is, $\mathcal{L}h_{i}=-\lambda_{i}h_{i}.$ The eigenvalues $\lambda_{i}$’s satisfy $\lambda_{1}<\lambda_{2}\leq\lambda_{3}\leq\cdots$ tending to +,$\infty$,+ ∞ , and each eigenspace is finite dimensional.
-The next thing we need is the nice decaying property of these eigenfunctions.
+The next thing we need is a nice decaying property of these eigenfunctions.
 Similar to the eigenfunctions of the weighted Lichnerowicz operator, one expects that the $L$-eigenfunctions also decay in a suitable manner.
-This part follows from a general result proven by Colding-Minicozzi in [CM21].
-A simpler case for $\mathcal{L}$-eigenfunctions was dealt with by Colding-Minicozzi in [CM20].
+This part follows from a general result proven by Colding–Minicozzi in [CM21].
+A simpler case for $\mathcal{L}$-eigenfunctions was dealt with by Colding–Minicozzi in [CM20].
 Theorem 2.4.
@@
 $h\mathcal{L}h\geq-\lambda|h|^{2}-\left(\delta+\frac{1}{2}\right)h^{2}.$
-Now we apply [CM21, Theorem 3.9] by considering $f=\frac{|x|^{2}}{4},$ $b=|x|=r,$ and $\psi=Ch^{2}$ in their work and obtain the desired estimate for $h.$ Note that the $\lambda$-eigenspace is finite-dimensional, so we can find a constant dependeing on $\lambda$ working for all $\lambda$-eigenfunctions.
+Now we apply [CM21, Theorem 3.9] by considering $f=\frac{|x|^{2}}{4},$ $b=|x|=r,$ and $\psi=Ch^{2}$ in their work and obtain the desired estimate for $h.$ Note that the $\lambda$-eigenspace is finite-dimensional, so we can find a constant depending on $\lambda$ working for all $\lambda$-eigenfunctions.
 ∎
@@
 The main idea is similar to [Sto], but here the process will be more technical since we are working on extrinsic hypersurfaces.
-As in (2.1), denoting the asymptotic cone of $\Sigma$ by $C$, from [W14, CS], we know that outside a large compact set, we can parameterize $\Sigma\setminus B_{R}(0)$ as
+As in (2.1), denoting the asymptotic cone of $\Sigma$ by $\mathcal{C}$, from [W14, CS], we know that outside a large compact set, we can parameterize $\Sigma\setminus B_{R}$ by
 $F_{0}(r,\theta)=r\theta+f(r,\theta)\textbf{n}$
-over $C$, where $\theta$ is the parametrization for the link of $C$ and n is the normal vector of $C$ at $(r,\theta)$.
+over $\mathcal{C}\setminus B_{R}$, where $\theta$ is the parametrization for the link of $C$ and n is the normal vector of $\mathcal{C}$ at $(r,\theta)$.
 We choose a smooth cutoff function $\eta\colon\mathbb{R}\to\mathbb{R}$ such that $\eta=\left\{\begin{array}[]{cc}1,\,\,x<0\\ 0,\,\,x>1\end{array}\right..$ Take $F_{1}(r,\theta)=r\theta+\eta(\frac{r-R}{R})f(r,\theta)\textbf{n}$.
@@
 The first step is to move the link to a hemisphere.
-As the link of $C$ is a codimension one submanifold of $S^{n-1}$, we can pick a vector $\mathbf{v}\in S^{n-1}$ such that $-\mathbf{v}$ is not contained in the link of $C$.
+As the link of $\mathcal{C}$ is a codimension one submanifold of $S^{n-1}$, we can pick a vector $\mathbf{v}\in S^{n-1}$ such that $-\mathbf{v}$ is not contained in the link of $\mathcal{C}$.
 Initially, we have $F_{1}(r,\theta)=r\theta$ for $r>2R.$ We take
@@
 This proposition seems very technical, but it enables us to easily control the $C^{0}$ norm of h using this barrier function.
-By combining Proposition 4, Lemma 4, Lemma 4, and Lemma 4, we can prove the global $C^{2}$-improvement theorem following the same argument in [Sto].
+By combining Proposition 4, Lemma 4, and Lemma 4, we can prove the global $C^{2}$-improvement theorem following the same argument in [Sto].
 Note that here we need $C^{1}$-bounds instead of $C^{0}$-bounds in the mean curvature flow setting.
@@
 $\|h_{\mathbf{p},s}(\cdot,-\log(1-t(\mathbf{p})))\|_{L^{2}_{W}}<\mu_{s}e^{\lambda_{*}(\tau_{0}-\log(1-t(\mathbf{p})))},$
-and thus, from the definition of the box, ww obtain
+and thus, from the definition of the box, we obtain
 $\|h_{\mathbf{p},u}(\cdot,-\log(1-t(\mathbf{p})))\|_{L^{2}_{W}}=\mu_{u}e^{\lambda_{*}(\tau_{0}-\log(1-t(\mathbf{p})))}.$
@@
 ∎
+Using Theorem 1 and the results from [AIV, CDS], we can prove Corollary 1.1.
+Proof of Corollary 1.1.
+Angenent–Ilmanen–Velázquez constructed $n$-dimensional asymptotically conical shrinkers such that level set flows starting from their asymptotic cones fatten for $n\in\left\{3,4,5,6\right\}$ in [AIV, Theorem 1].
+[^1] The proof of Theorem 1 in [AIV] was provided in Section 3 (for proving Theorem 5) and Section 5.1 there.
+Take such a shrinker $\Sigma$ and let $\mathcal{C}$ be its asymptotic cone.
+Theorem 1 implies that there exists a smooth closed hypersurface $M\subseteq\mathbb{R}^{n+1}$ such that the level set flow $M_{t}$ starting from $M$ develops a singularity at $({\bf 0},0)\in\mathbb{R}^{n+1}\times\mathbb{R}$ whose tangent flow is modeled on $\Sigma.$ Thus, $M_{0}$ is a hypersurface with a singularity at the origin modeled on the cone $C$ (cf.
+[CS, Corollary 1.2]).
+Since the level set flow starting from $\mathcal{C}$ fattens, the main result in [CDS] implies that the level set flow $M_{t}$ fattens for positive time.
+∎

### v3

@@
 As a corollary, our construction, combined with the works of Angenent–Ilmanen–Velázquez and Chodosh–Daniels-Holgate–Schulze, implies the existence of fattening level set flows starting from smooth embedded closed hypersurfaces.
-This addresses a question posed by Evans–Spruck and De Giorgi.
+These provide examples related to a question asked by Evans–Spruck.
 1. Introduction
@@
 He described the proof ideas in [W02, Section 4], which are completely different from our approach to Corollary 1.1.
-Combining their result with Corollary 1.1, we now have $n$-dimensional embedded closed examples for $n\in\left\{2,3,4,5,6\right\}.$
+In a recent work [IW], Ilmanen–White improved the result and showed that for any large enough $g,$ there exists a smooth embedded closed surface in $\mathbb{R}^{3}$ of genus $g$ such that the level set flow starting from it develops an asymptotically conical singularity (with two ends) and fattens after that singular time.
+Their work was based on the resolution of the multiplicity one conjecture by Bamler–Kleiner [BK23].
+Combining their results with Corollary 1.1, we now have $n$-dimensional embedded closed fattening examples for $n\in\left\{2,3,4,5,6\right\}.$
+There was another recent work by Ketover [K24], where he used min-max methods to construct a genus $g$ asymptotically conical shrinker with two ends and an entropy estimate.
+Furthermore, the level set flow starting from the asymptotic cone of the shrinker fattens when $g$ is large.
+As pointed out in [IW], combining Ketover’s result with Theorem 1 and arguing as in the proof of Corollary 1.1 leads to other examples of two-dimensional fattening level set flows starting from smooth embedded closed surfaces.
+We briefly mention the strategy to prove Theorem 1.
 The proof of Theorem 1 follows the outline of the argument of the corresponding construction in the Ricci flow case by Stolarski [Sto].
@@
 The goal is to make the mean curvature flow starting from a perturbation of a doubling of an asymptotically conical shrinker develop a singularity modeled on the chosen shrinker itself.
-There are three main difficulties in the case of mean curvature flows.
+To be more specific, we perturb the doubling by eigenfunctions and consider the mean curvature flow starting from the perturbations.
+By considering the first time when the flow gets outside of the ”boxes with good estimates”, we can see that the only possibility is that its unstable component hits the boundary.
+By topological arguments, we can prove that there exists a perturbation which always stays in the box, which means that it develops the corresponding singularity.
+An obvious difference between our work and [Sto] is the extrinsic nature of mean curvature flows.
+When working with the Ricci flow, Stolarski perturbed the metric tensor to achieve the goal of constructing the desired closed Riemannian manifold.
+In the case of mean curvature flows, we perturb a doubling of a shrinker by, as expected, looking at normal graphs on it.
+Thus, the role of the perturbing function $h$ in our work is not the role of the perturbing tensor in [Sto].
+Instead, since metric tensors come from pairings of first derivatives, perturbing tensors should correspond to the gradient of our perturbing functions.
+This requires additional work to relate the usual $C^{k}$ norms with the homogeneous norms on an asymptotically conical shrinker.
+Other than a slight difference mentioned above, there are three main difficulties in the case of mean curvature flows.
 The first one is that we have to write the flow as a normal graph over the shrinker to use the tools from analysis, while in the Ricci flow case [Sto], one only needs to pull back tensors.
-The second difficulty is that we have to guarantee the embeddedness of the modified doubling of an asymptotically conical shrinker and control the curvature bound on the gluing part, while in the Ricci flow case, we can use cutoff functions to modify the metric directly.
-The third difficulty is that the shrinker quantity is more nonlinear than the Ricci curvature and hence complicates the arguments.
+Moreover, we have to guarantee the embeddedness of the modified doubling of an asymptotically conical shrinker and control the curvature bound on the gluing part, while in the Ricci flow case, we can use cutoff functions to modify the metric directly.
+Finally, the shrinker quantity is more nonlinear than the Ricci curvature and hence complicates the arguments.
+We remark that there are constructions of mean curvature flows with singularities modeled on minimal cones, which form another important class of singularities.
+In [V94], Velázquez constructed mean curvature flows with singularities modeled on the Simons cones.
+Following his work, more precise information has been studied for flows with such singularities.
+See, for example, [GS18, S23V, Liu].
 This paper is organized as follows.
@@
 In Sections 4 and 5, we talk about the interior estimates and constructions of barrier functions to get the preservation and the improvement of $C^{2}$ bounds.
-In Section 6, we use the Ważewski box argument to finish the proof of Theorem 1 and prove Corollary 1.1.
+In Section 6, we use the Ważewski box argument to finish the proof of Theorem 1 and conclude with the proof of Corollary 1.1.
 Acknowledgement
-The authors are grateful to Prof.
-Toby Colding and Prof.
-Bill Minicozzi for their constant support and several inspiring conversations.
-They would like to thank Letian Chen, Otis Chodosh, Alec Payne, Max Stolarski, and Ao Sun for valuable discussions, and to Alec Payne for bringing the work of Angenent, Ilmanen, and Velázquez to their attention and suggesting the application regarding fattening.
-During the project, Lee was partially supported by NSF Grant DMS 2005345 and Zhao was supported by NSF Grant DMS 1812142 and NSF Grant DMS 1811267.
+The authors are grateful to Toby Colding and Bill Minicozzi for their constant support and several inspiring conversations.
+They would like to thank Letian Chen, Otis Chodosh, Alec Payne, Max Stolarski, Ao Sun, Lu Wang, and Brian White for valuable discussions and comments, and to Alec Payne for bringing the work of Angenent, Ilmanen, and Velázquez to their attention and suggesting the application regarding fattening.
+During the project, Lee was partially supported by NSF Grant DMS 2005345 and Zhao was supported by NSF Grant DMS 1812142, NSF Grant DMS 1811267 and NSF Grant DMS 2104349.
 2. Preliminaries facts and evolution equations
-2.1. Stability operators on asymptotically conical shrinkers
+2.1. Analysis on asymptotically conical shrinkers
 Let $\Sigma$ be an asymptotically conical shrinker in $\mathbb{R}^{n+1}.$ In this paper, this means that its asymptotic cone
@@
 Proposition 2.3.
-There exists an $L^{2}_{W}$-orthonormal basis $\{h_{i}\}_{i\in\mathbb{N}}$ of $L^{2}_{W}(\Sigma)$ such that for all $i\in\mathbb{N},$ $h_{i}$ is an eigenfunction of $\mathcal{L}$ with eigenvalue $\lambda_{i},$ that is, $\mathcal{L}h_{i}=-\lambda_{i}h_{i}.$ The eigenvalues $\lambda_{i}$’s satisfy $\lambda_{1}<\lambda_{2}\leq\lambda_{3}\leq\cdots$ tending to +,$\infty$,+ ∞ , and each eigenspace is finite dimensional.
+There exists an $L^{2}_{W}$-orthonormal basis $\{h_{i}\}_{i\in\mathbb{N}}$ of $L^{2}_{W}(\Sigma)$ such that for all $i\in\mathbb{N},$ $h_{i}$ is an eigenfunction of $L$ with eigenvalue $\lambda_{i},$ that is, $Lh_{i}=-\lambda_{i}h_{i}.$111 We use the different sign convention about eigenvalues from the one used in [Sto].
+The eigenvalues $\lambda_{i}$’s satisfy $\lambda_{1}<\lambda_{2}\leq\lambda_{3}\leq\cdots$ tending to +,$\infty$,+ ∞ , and each eigenspace is finite dimensional.
 The next thing we need is a nice decaying property of these eigenfunctions.
@@
 If $Lh=-\lambda h$ with $\|h\|_{L^{2}_{W}}=1,$ then
-$|h(x)|\leq C|x|^{\lambda+\delta+\frac{1}{2}}.$
+$|h(x)|\leq C|x|^{2\left(\lambda+\delta+\frac{1}{2}\right)}.$
 Proof.
@@
 ∎
+Finally, we mention some weighted spaces that we will use following [CS].
+Let $\tilde{r}$ be a fixed smooth function such that $\tilde{r}\geq 1$ and $\tilde{r}=|x|$ for $|x|\geq 2$.
+We consider the weighted norms
+$\|f\|_{0;-\gamma}^{\rm hom}$ $:=\sup_{x\in\Sigma}\tilde{r}(x)^{\gamma}|f(x)|,\text{and}$
+$[f]_{\alpha;-\gamma-\alpha}^{\rm hom}$ $:=\sup_{x,y\in\Sigma}\frac{1}{\tilde{r}(x)^{-\gamma-\alpha}+\tilde{r}(y)^{-\gamma-\alpha}}\frac{|f(x)-f(y)|}{|x-y|^{\alpha}}.$
+The space $C^{0,\alpha}_{{\rm hom},-\gamma}$ is defined to collect all $f$ such that
+$\|f\|_{C^{0,\alpha}_{{\rm hom},-\gamma}}:=\|f\|_{0;-\gamma}^{\rm hom}+[f]^{\rm hom}_{\alpha;-\gamma-\alpha}<\infty,$
+and the space $C^{1,\alpha}_{{\rm hom},-\gamma}$ is defined to collect all $f$ such that
+$\|f\|_{C^{1,\alpha}_{{\rm hom},-\gamma}}:=\sum_{j=0}^{1}\|(\nabla_{\Sigma})^{j}f\|_{C^{0,\alpha}_{{\rm hom},-j-\gamma}}<\infty.$
 2.2. Evolution equation of immersions
@@
 so $\varphi_{t}(x)$ satisfies
-$(\partial_{t}F_{0})(\varphi_{t}(x),t)+dF_{0}(\partial_{t}\varphi_{t}(x))(\varphi_{t}(x),t)$
-$=$ $\sqrt{1-t}\,\partial_{t}{h}(x,t)\textbf{n}_{F}(x)-\frac{1}{2\sqrt{1-t}}\left(F(x)+{h}(x,t)\textbf{n}_{F}(x)\right).$
+$(\partial_{t}F_{0})(\varphi_{t}(x),t)+dF_{0}(\partial_{t}\varphi_{t}(x))(\varphi_{t}(x),t)=\sqrt{1-t}\,\partial_{t}{h}(x,t)\textbf{n}_{F}(x)-\frac{1}{2\sqrt{1-t}}\left(F(x)+{h}(x,t)\textbf{n}_{F}(x)\right).$
 Thus, we have
-$(\partial_{t}F_{0})(\varphi_{t}(x),t)+dF_{0}(\partial_{t}\varphi_{t}(x))(\varphi_{t}(x),t)+\frac{1}{2\sqrt{1-t}}F(x)$
-$=$ $(\sqrt{1-t}\,\partial_{t}{h}(x,t)-\frac{1}{2\sqrt{1-t}}{h}(x,t))\textbf{n}_{F}(x).$
-As we assume that locally $\frac{1}{\sqrt{1-t}}F_{1}$ can be written as a graph over $F$, we have
-$\langle\textbf{n}_{F_{0}}(\varphi_{t}(x)),\textbf{n}_{F}(x)\rangle\neq 0.$
-We denote
+$(\partial_{t}F_{0})(\varphi_{t}(x),t)+dF_{0}(\partial_{t}\varphi_{t}(x))(\varphi_{t}(x),t)+\frac{1}{2\sqrt{1-t}}F(x)=\left(\sqrt{1-t}\,\partial_{t}{h}(x,t)-\frac{1}{2\sqrt{1-t}}{h}(x,t)\right)\textbf{n}_{F}(x).$
+As we assume that locally $\frac{1}{\sqrt{1-t}}F_{1}$ can be written as a graph over $F$, we have $\langle\textbf{n}_{F_{0}}(\varphi_{t}(x)),\textbf{n}_{F}(x)\rangle\neq 0.$ We write
 (2.6) $\textbf{n}_{F}(x)=\langle\textbf{n}_{F}(x),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle\textbf{n}_{F_{0}}(\varphi_{t}(x))+(\textbf{n}_{F}(x)-\langle\textbf{n}_{F}(x),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle\textbf{n}_{F_{0}}(\varphi_{t}(x))),$
@@
 In particular, we have
-(2.7) $|dF_{0}(\partial_{t}\varphi_{t}(x))|\leq\frac{\left|(\partial_{t}F_{0})(\varphi_{t}(x),t)+\frac{1}{2\sqrt{1-t}}F(x)\right|}{|\langle\textbf{n}_{F}(x),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle|}.$
+$|dF_{0}(\partial_{t}\varphi_{t}(x))|\leq\frac{\left|(\partial_{t}F_{0})(\varphi_{t}(x),t)+\frac{1}{2\sqrt{1-t}}F(x)\right|}{|\langle\textbf{n}_{F}(x),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle|}.$
 Therefore, we have
 (2.7) $\partial_{t}{h}(x,t)=\frac{1}{2(1-t)}{h}(x,t)+\frac{\frac{1}{\sqrt{1-t}}\langle(\partial_{t}F_{0})(\varphi_{t}(x),t),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle+\frac{1}{2(1-t)}\langle F(x),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle}{\langle\textbf{n}_{F}(x),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle}.$
-Setting $e^{\tau}=\frac{1}{1-t},$ we get
-$\partial_{t}=\frac{\partial\tau}{\partial t}\partial_{\tau}=\frac{1}{1-t}\partial_{\tau},$
-and thus,
+Setting $e^{\tau}=\frac{1}{1-t},$ we get $\partial_{t}=\frac{\partial\tau}{\partial t}\partial_{\tau}=\frac{1}{1-t}\partial_{\tau},$ and thus,
 $\partial_{\tau}{h}(x,t)=$ $\frac{1}{2}{h}(x,t)+\frac{{\sqrt{1-t}}\langle(\partial_{t}F_{0})(\varphi_{t}(x),t),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle+\frac{1}{2}\langle F(x),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle}{\langle\textbf{n}_{F}(x),\textbf{n}_{F_{0}}(\varphi_{t}(x))\rangle}$
@@
 (2.8) $\widetilde{\textbf{n}}$ $=-\nabla^{g}{h}+C_{1}(F,\nabla F,\nabla^{2}F,h,\nabla h)*\nabla(h\,\textbf{n}_{F})*\nabla(h\,\textbf{n}_{F}),$
-$\hat{h}_{ab}$ $=\langle\partial_{a}F,\partial_{b}(h\,\textbf{n}_{F})\rangle+\langle\partial_{a}(h\,\textbf{n}_{F}),\partial_{b}F\rangle+\langle\partial_{a}(h\,\textbf{n}_{F}),\partial_{b}(h\,\textbf{n}_{F})\rangle=2hA_{ab}+\langle\partial_{a}h,\partial_{b}h\rangle+h^{2}A_{ia}A_{bj}g^{ij},$
+$\hat{h}_{ab}$ $=\langle\partial_{a}F,\partial_{b}(h\,\textbf{n}_{F})\rangle+\langle\partial_{a}(h\,\textbf{n}_{F}),\partial_{b}F\rangle+\langle\partial_{a}(h\,\textbf{n}_{F}),\partial_{b}(h\,\textbf{n}_{F})\rangle=2hA_{ab}+\langle\partial_{a}h,\partial_{b}h\rangle+h^{2}A_{ia}A_{bj}g^{ij},\text{and}$
 $\tilde{h}^{ab}$ $=-2hA^{ab}+C_{2,ab}(F,\nabla F,\nabla^{2}F,h,\nabla h)*\nabla h*\nabla h+C_{3,ab}(F,\nabla F,\nabla^{2}F,h,\nabla h)*hA,$
 where
-$\|C_{1}(F,\nabla F,\nabla^{2}F,h,\nabla h)\|_{C^{0}}+\sum_{a,b}\|C_{2,ab}(F,\nabla F,\nabla^{2}F,h,\nabla h)\|_{C^{0}}+\|C_{3,ab}(F,\nabla F,\nabla^{2}F,h,\nabla h)\|_{C^{0}}\leq C_{0}\|h\|_{C^{1}}.$
+$\|C_{1}(F,\nabla F,\nabla^{2}F,h,\nabla h)\|_{C^{0}}+\sum_{a,b}\|C_{2,ab}(F,\nabla F,\nabla^{2}F,h,\nabla h)\|_{C^{0}}+\|C_{3,ab}(F,\nabla F,\nabla^{2}F,h,\nabla h)\|_{C^{0}}\leq C_{0}\|h\|_{C^{1,\alpha}_{\rm hom,-1}}.$
 Also, note that $\Delta^{g}\textbf{n}_{F}=-|A|^{2}\textbf{n}_{F}+\nabla^{g}H$ and $A_{ij,k}=A_{ik,j}$ in $\mathbb{R}^{n+1}$.
-Here, we use the notation $\textbf{H}=\Delta^{g}F$ and $H=-\textbf{H},\textbf{n}_{F}\rangle=g^{ij}A_{ij},$ so in normal coordinates, we have
+Here, we use the notation $\textbf{H}=\Delta^{g}F$ and $H=g^{ij}A_{ij}=-\left\langle\textbf{H},\textbf{n}_{F}\right\rangle,$ so in normal coordinates, we have
 $\left\langle\Delta^{g}h\cdot\textbf{n}_{F}-h|A|^{2}\textbf{n}_{F}+\frac{1}{2}h\textbf{n}_{F}+\frac{1}{2}(\langle F,\widetilde{\textbf{n}}\rangle\textbf{n}_{F}+\langle F,\textbf{n}_{F}\rangle\widetilde{\textbf{n}})+\tilde{h}^{ab}\nabla^{g}_{a}\nabla^{g}_{b}F,\textbf{n}_{F}\right\rangle$
@@
 The main idea is similar to [Sto], but here the process will be more technical since we are working on extrinsic hypersurfaces.
-As in (2.1), denoting the asymptotic cone of $\Sigma$ by $\mathcal{C}$, from [W14, CS], we know that outside a large compact set, we can parameterize $\Sigma\setminus B_{R}$ by
+As in (2.1), denoting the asymptotic cone of $\Sigma$ by $\mathcal{C}$, from [W14, CS], we know that outside a large compact set $\Sigma\cap\overline{B}_{R}$, we can parameterize $\Sigma\setminus B_{R}$ by
 $F_{0}(r,\theta)=r\theta+f(r,\theta)\textbf{n}$
-over $\mathcal{C}\setminus B_{R}$, where $\theta$ is the parametrization for the link of $C$ and n is the normal vector of $\mathcal{C}$ at $(r,\theta)$.
+over $\mathcal{C}\setminus B_{R}$, where $\theta$ is the parametrization for the link of $\mathcal{C}$ and n is the normal vector of $\mathcal{C}$ at $(r,\theta)$.
+We will modify $\Sigma$ in this region, and glue it with the compact part.
 We choose a smooth cutoff function $\eta\colon\mathbb{R}\to\mathbb{R}$ such that $\eta=\left\{\begin{array}[]{cc}1,\,\,x<0\\ 0,\,\,x>1\end{array}\right..$ Take $F_{1}(r,\theta)=r\theta+\eta(\frac{r-R}{R})f(r,\theta)\textbf{n}$.
@@
 so that for $r\geq 3R,$ the image is contained in a hemisphere. Consequently, we can define
-$F_{3}(r,\theta)=\begin{cases}r\theta+\eta(\frac{r-R}{R})f(r,\theta)\textbf{n}&\text{if }r\leq 2R\\ r\cdot\frac{\left(\frac{2}{3}\eta\left(\frac{r-2R}{R}\right)+\frac{1}{3}\right)\theta+\left(\frac{2}{3}-\frac{2}{3}\eta\left(\frac{r-2R}{R}\right)\right)\mathbf{v}}{\left|\left(\frac{2}{3}\eta\left(\frac{r-2R}{R}\right)+\frac{1}{3}\right)\theta+\left(\frac{2}{3}-\frac{2}{3}\eta\left(\frac{r-2R}{R}\right)\right)\mathbf{v}\right|}&\text{if }r\geq 2R\end{cases}.$
+$F_{3}(r,\theta)=\begin{cases}r\theta+\eta(\frac{r-R}{R})f(r,\theta)\textbf{n}&\text{if }R\leq r\leq 2R\\ r\cdot\frac{\left(\frac{2}{3}\eta\left(\frac{r-2R}{R}\right)+\frac{1}{3}\right)\theta+\left(\frac{2}{3}-\frac{2}{3}\eta\left(\frac{r-2R}{R}\right)\right)\mathbf{v}}{\left|\left(\frac{2}{3}\eta\left(\frac{r-2R}{R}\right)+\frac{1}{3}\right)\theta+\left(\frac{2}{3}-\frac{2}{3}\eta\left(\frac{r-2R}{R}\right)\right)\mathbf{v}\right|}&\text{if }r\geq 2R\end{cases}.$
 The next step is to modify the end such that the radial derivative is in the same direction. To this end, we take
@@
 […]
+We can now glue $F$ with the original shrinker along $F(R,\cdot)$ with $\Sigma\cap\overline{B}_{R}$.
+As $\eta$ is a cutoff function, we have that the glued hypersurface is also smooth.
 From the constructions, the $F_{3}$ part lies in a hemisphere and controls the distance to origin.
@@
 Given $\Gamma>0$ and $h$ satisfying (2.9) on $\Sigma\times(0,\tau_{1})$, assume that $\mathcal{Q}_{1}=0$ on $\left\{\frac{|F|^{2}}{4}<2\Gamma e^{\tau_{0}}\right\}\times[0,\tau_{1}]$ and
-$\|h\|_{L^{\infty}(\Sigma\times[0,\tau_{1}])}$ $<\varepsilon,$
-$\|h(\tau)\|_{L_{\frac{|f|^{2}}{4}}^{2}(\Sigma)}$ $\leq\mu e^{\lambda_{*}(\tau+\tau_{0})}\,\,\text{ for all}\,\,\tau\in[0,\tau_{1}],\text{ and}$
-$\sum_{l=0}^{4n+3}\left|\nabla^{l}h\right|_{g}(x,0)$ $\leq C_{0}\bar{p}e^{\lambda_{*}\tau_{0}}\,\,\text{ for all}\,\,x\in\left\{\frac{|F|^{2}}{4}<2\Gamma e^{\tau_{0}}\right\}\text{ and }\tau=0.$
-Then for $\Gamma>\Gamma_{0}(n,\Sigma,F)$, $\varepsilon<\varepsilon_{0}(n,\Sigma,F,\Gamma),$ and $\tau_{0}>C(n,\Sigma,F,\lambda_{*},\Gamma,C_{0}\bar{p})$, there exists $C_{1}=C_{1}(n,\Sigma,F,\lambda_{*},\Gamma)<\infty$ such that
-$\sum_{l=0}^{2}|\nabla^{g}h|_{g}(x,\tau)\leq C_{1}(\mu+C_{0}\bar{p})e^{\lambda_{*}(\tau+\tau_{0})}\,\,\text{ for all}\,\,(x,\tau)\in\left\{\frac{|F|^{2}}{4}<\Gamma e^{\tau_{0}}\right\}\times[0,\tau_{1}].$
+$\|h\|_{C^{0,\alpha}_{\rm hom,-1}(\Sigma\times[0,\tau_{1}])}$ $<\varepsilon,$
+$\|h(\tau)\|_{L^{2}_{W}(\Sigma)}$ $\leq\mu e^{-\lambda_{*}(\tau+\tau_{0})}\,\,\text{ for all}\,\,\tau\in[0,\tau_{1}],\text{ and}$
+$\sum_{l=0}^{4n+3}\left|\nabla^{l}h\right|_{g}(x,0)$ $\leq C_{0}\bar{p}e^{-\lambda_{*}\tau_{0}}\,\,\text{ for all}\,\,x\in\left\{\frac{|F|^{2}}{4}<2\Gamma e^{\tau_{0}}\right\}\text{ and }\tau=0$
+for some $\lambda_{*}>0.$ Then for $\Gamma>\Gamma_{0}(n,\Sigma,F)$, $\varepsilon<\varepsilon_{0}(n,\Sigma,F,\Gamma),$ and $\tau_{0}>C(n,\Sigma,F,\lambda_{*},\Gamma,C_{0}\bar{p})$, there exists $C_{1}=C_{1}(n,\Sigma,F,\lambda_{*},\Gamma)<\infty$ such that
+$\sum_{l=0}^{2}|\nabla^{g}h|_{g}(x,\tau)\leq C_{1}(\mu+C_{0}\bar{p})e^{-\lambda_{*}(\tau+\tau_{0})}\,\,\text{ for all}\,\,(x,\tau)\in\left\{\frac{|F|^{2}}{4}<\Gamma e^{\tau_{0}}\right\}\times[0,\tau_{1}].$
 Note that we can also use Proposition 4 to get large scale interior estimates by rescaling. More precisely, consider a family of diffeomorphisms $\varphi_{s}\colon M\to M$ given by
@@
 This helps us to get rid of the unbounded coefficient term $-\frac{1}{2}\langle F,\nabla^{g}f\rangle$ in $L$.
-Also, note that $0\leq\left|\nabla^{g}(\frac{|F|^{2}}{4})\right|^{2}\leq\frac{|F|^{2}}{4}$ gives us that for $s_{1}<s_{2},$
+Also, note that the relation $0\leq\left|\nabla^{g}(\frac{|F|^{2}}{4})\right|^{2}\leq\frac{|F|^{2}}{4}$ implies that for $s_{1}<s_{2},$
 $\frac{|F|^{2}}{4}(\varphi_{s_{1}}(x))\leq\frac{|F|^{2}}{4}(\varphi_{s_{2}}(x))$
@@
 $e^{-s_{1}}\frac{|F|^{2}}{4}(\varphi_{s_{1}}(x))\geq e^{-s_{2}}\frac{|F|^{2}}{4}(\varphi_{s_{2}}(x)).$
-Thus, Proposition 4 implies the large scale interior estimates.
+Moreover, from [CS, Lemma 3.8], if we denote the link of the asymptotic cone of $\Sigma$ by $\mathcal{C}$, by parameterizing the shrinker outside a compact set by $F:(R,\infty)\times\Gamma\to\Sigma$ and denoting the family of diffeomorphisms generated by $X_{t}=\frac{1}{2(-t)}x^{T}$ by $\Phi_{t}$, we can define $\tilde{\Phi}_{t}=F^{-1}\circ\Phi_{t}\circ F$ and $\phi_{t}:(R,\infty)\times\Gamma\to(R,\infty)\times\Gamma,\,(r,\omega)\to((-t)^{-\frac{1}{2}}r,\omega)$ with the following estimates.
+[…]
+$d_{g_{\mathcal{C}}}(\tilde{\Phi_{t}}(t,\theta),\phi_{t}(r,\theta))\lesssim\frac{1}{\sqrt{-t}\cdot r},$
+$|D^{(j)}(\tilde{\Phi}-\phi_{t})|(r,\theta)\lesssim\frac{1}{\sqrt{-t}\cdot r^{1+j-\eta}},$
+for $j\geq 1$ and $\eta>0$.
+As in [CS], we recall a non-standard interior Schauder estimates from [Bra69].
+[…]
+Suppose $B_{2}\in\mathbb{R}^{n}$ and there are $C^{0,\alpha}$ functions $a_{ij},b_{i},c,u,f\colon B_{2}\times[-2,0]\to\mathbb{R}$ such that
+$\frac{\partial}{\partial t}u=a_{ij}D_{ij}^{2}u+b_{i}D_{i}u+cu+f.$
+If the system is uniformly parabolic in the sense that $a_{ij}\xi_{i}\xi_{j}\geq\lambda|\xi|^{2}$ for some $\lambda>0$ and
+$\sup_{t\in[-2,0]}\left(\|a_{ij}(\cdot,t)\|_{C^{0,\alpha}(B_{1})}+\|b_{i}(\cdot,t)\|_{C^{0,\alpha}(B_{1})}+\|c(\cdot,t)\|_{C^{0,\alpha}(B_{1})}\right)\leq\Lambda$
+for some $\Lambda>0,$ then for $T\in(-1,0],$
+$\sup_{t\in[-1,T]}\|u(\cdot,t)\|_{C^{2,\alpha}(B_{1})}\leq C\sup_{t\in[-2,T]}\left(\|u(\cdot,t)\|_{C^{0,\alpha}(B_{2})}+\|f(\cdot,t)\|_{C^{0,\alpha}(B_{2})}\right)$
+for some $C=C(n,\alpha,\lambda,\Lambda)$.
+Combining Lemma 4, Lemma 4 and Proposition 4 implies the large scale interior estimates.
 […]
 Given $h$ satisfying (2.9) and $\mathcal{Q}_{1}=0$ on $\left\{(x,\tau)\in\Sigma\times(0,\tau_{1}):\,\Gamma e^{\tau_{0}}<\frac{|F|^{2}}{4}<\gamma e^{\tau+\tau_{0}}\right\},$ assume that
-$|h|_{g}$ $\leq C_{0}e^{-k(\tau+\tau_{0})}\left(\frac{|F|^{2}}{4}(x)\right)^{k}\,\,\text{on}\,\,\left\{(x,\tau)\in\Sigma\times(0,\tau_{1}):\,\Gamma e^{\tau_{0}}<\frac{|F|^{2}}{4}<\gamma e^{\tau+\tau_{0}}\right\},\text{ and}$
-$\sum_{l=0}^{3}|\nabla^{l}h|_{g}(x,\tau_{0})$ $\leq C_{0}e^{-k\tau_{0}}\left(\frac{|F|^{2}}{4}(x)\right)^{k}\,\,\text{for all}\,\,x\in\left\{\Gamma e^{\tau_{0}}<\frac{|F|^{2}}{4}<\gamma e^{\tau+\tau_{0}}\right\}.$
+$|h|_{g}$ $\leq C_{0}e^{-k(\tau+\tau_{0})}\left(\frac{|F|^{2}}{4}(x)\right)^{k+\frac{1}{2}}\,\,\text{on}\,\,\left\{(x,\tau)\in\Sigma\times(0,\tau_{1}):\,\Gamma e^{\tau_{0}}<\frac{|F|^{2}}{4}<\gamma e^{\tau+\tau_{0}}\right\},\text{ and}$
+$|\nabla^{l}h|_{g}(x,\tau_{0})$ $\leq C_{0}e^{-k\tau_{0}}\left(\frac{|F|^{2}}{4}(x)\right)^{k+\frac{1}{2}-\frac{l}{2}}\,\,\text{for all}\,\,l\in\{0,1,2,3\},\,\,x\in\left\{\Gamma e^{\tau_{0}}<\frac{|F|^{2}}{4}<\gamma e^{\tau+\tau_{0}}\right\}.$
 Then for $\Gamma>\Gamma_{0}(n,\Sigma,F)$ and $C_{0}\gamma^{k}<C(n,\Sigma,F)$, there exists $C_{1}=C_{1}(n,\Sigma,F,k)<\infty$ such that
-$\sum_{l=0}^{2}|\nabla^{g}h|_{g}(x,\tau)\leq C_{1}C_{0}e^{-k(\tau+\tau_{0})}\left(\frac{|F|^{2}}{4}(x)\right)^{k}\,\,\text{on}\,\,\left\{(x,\tau)\in\Sigma\times[0,\tau_{1}]:\,4\Gamma e^{\tau_{0}}\leq\frac{|F|^{2}}{4}(x)\leq\frac{1}{4}\gamma e^{\tau+\tau_{0}}\right\}.$
-Similarly, on the outer region, we have the following estimate.
+$|\nabla^{l}h|_{g}(x,\tau)\leq C_{1}C_{0}e^{-k(\tau+\tau_{0})}\left(\frac{|F|^{2}}{4}(x)\right)^{k+\frac{1}{2}-\frac{l}{2}}\,\,\text{on}\,\,\left\{(x,\tau)\in\Sigma\times[0,\tau_{1}]:\,4\Gamma e^{\tau_{0}}\leq\frac{|F|^{2}}{4}(x)\leq\frac{1}{4}\gamma e^{\tau+\tau_{0}}\right\}$
+for all $l\in\{0,1,2\}.$
+Similarly, in the outer region, we have the following estimate.
 […]
 Given $h$ satisfying (2.9) on $\Sigma\times(0,\tau_{1}),$ assume that
-$|h|_{g}(x,\tau)$ $\leq\delta\,\,\text{on}\,\,\left\{(x,t)\in\Sigma\times[0,\tau_{1}]:\,\gamma e^{\tau+\tau_{0}}<\frac{|F|^{2}}{4}(x)\right\},$
-$\sum_{l=0}^{3}|\nabla^{l}h|_{g}(x,0)$ $\leq\delta\,\,\text{on}\,\,x\in\left\{\gamma e^{\tau_{0}}\leq\frac{|F|^{2}}{4}\right\},\text{ and}$
+$|h|_{g}(x,\tau)$ $\leq\delta|F|\,\,\text{on}\,\,\left\{(x,t)\in\Sigma\times[0,\tau_{1}]:\,\gamma e^{\tau+\tau_{0}}<\frac{|F|^{2}}{4}(x)\right\},$
+$|\nabla^{l}h|_{g}(x,0)$ $\leq\delta|F|^{1-l},\,\,l\in\{0,1,2,3\}\,\,\text{on}\,\,x\in\left\{\gamma e^{\tau_{0}}\leq\frac{|F|^{2}}{4}\right\},\text{ and}$
 $|\mathcal{Q}_{1}|_{g}(x,\tau))+|\nabla^{g}\mathcal{Q}_{1}|_{g}(x,\tau)$ $\leq\delta\,\,\text{on}\,\,\left\{(x,\tau)\in\Sigma\times[0,\tau_{1}]:\,\gamma e^{\tau+\tau_{0}}<\frac{|F|^{2}}{4}(x)\right\}.$
 Then for $\delta<\delta_{0}(n,\Sigma,F)$ and $\tau_{0}>\tau(n,\Sigma,F,\gamma)$, there exists $C=C(n,\Sigma,F)<\infty$ such that
-$\sum_{l=0}^{2}|\nabla^{g}h|_{g}(x,\tau)\leq C\delta\,\,\text{on}\,\,\left\{(x,\tau)\in\Sigma\times[0,\tau_{1}]:\,4\gamma e^{\tau+\tau_{0}}<\frac{|F|^{2}}{4}(x)\right\}.$
+$|\nabla^{l}h|_{g}(x,\tau)\leq C\delta|F|^{1-l},l\in\{0,1,2\}\,\,\text{on}\,\,\left\{(x,\tau)\in\Sigma\times[0,\tau_{1}]:\,4\gamma e^{\tau+\tau_{0}}<\frac{|F|^{2}}{4}(x)\right\}.$
 5. Barrier functions
@@
 We look at the lower order terms of $\mathcal{Q}_{2}$ in (2.9).
-Using (2.9), if we denote $C=C(F,\nabla F,\nabla^{2}F,h,\nabla h)$ to be any tensor depending on $F,\nabla F,\nabla^{2}F,h,\nabla h$, we get
+Using (2.8), if we denote $C=C(F,\nabla F,\nabla^{2}F,h,\nabla h)$ to be any tensor depending on $F,\nabla F,\nabla^{2}F,h,$ and $\nabla h$, we get
 $C_{3}h+C_{4}*\nabla h$
@@
 $\quad+2h\mathcal{Q}_{1}+h\left(C|\nabla h|^{2}+C|h||A|^{2}+C|h||\nabla A|\right).$
-Using (5.1), given $|h|_{C_{1}}\leq\varepsilon<\varepsilon_{n},$ we have
+Using (5.1), given $\|h\|_{C^{1,\alpha}_{\rm hom,-1}}\leq\varepsilon<\varepsilon_{n},$ we have
 $|C^{ab}|\leq C_{n}\left(|hA|_{C^{0}}+|\nabla h|^{2}_{C^{0}}\right),$
@@
 similarly as in [Sto], we take
-$u=\begin{cases}e^{(-k+1)(\tau+\tau_{0})}\left(Df^{k}-Bf^{k-1}\right)&\text{on }\left\{(x,\tau)\in\Sigma\times[0,\infty):\Gamma<f(x)<\gamma_{1}e^{\tau+\tau_{0}}\right\}\\ e^{\tau+\tau_{0}}\left(a-Bf^{-1}\right)&\text{on }\left\{(x,\tau)\in\Sigma\times[0,\infty):\gamma_{2}e^{\tau+\tau_{0}}<f(x)<\Gamma e^{\tau+\tau_{0}}\right\}\end{cases}.$
+$u=\begin{cases}e^{(-k+1)(\tau+\tau_{0})}\left(Df^{k}-Bf^{k-1}\right)&\text{on }\left\{(x,\tau)\in\Sigma\times[0,\infty):\Gamma<f(x)<\gamma_{1}e^{\tau+\tau_{0}}\right\}\\ e^{\tau+\tau_{0}}\left(a-Ef^{-1}\right)&\text{on }\left\{(x,\tau)\in\Sigma\times[0,\infty):\gamma_{2}e^{\tau+\tau_{0}}<f(x)<\Gamma e^{\tau+\tau_{0}}\right\}\end{cases}$
+for some positive numbers $a,B,D,$ and $E.$
 In general, using (5.3) and $0\leq\frac{|\nabla f|^{2}}{f}\leq 1,$ we calculate
@@
 Therefore, as $\mathcal{Q}_{1}$ is 0 on $\left\{(x,\tau)\in\Sigma\times[0,\infty):\Gamma e^{\tau_{0}}<f(x)<\gamma_{1}e^{\tau+\tau_{0}}\right\}$, there exist constants $\varepsilon_{n},\Gamma,$ and $\gamma_{1}$ such that given
-$|h|_{C_{1}}\leq\varepsilon<\varepsilon_{n},\quad B-D\left(m(m+\frac{n}{2})+C_{n}\sup_{\Sigma}\left(f|A|^{2}+f|\nabla A|\right)\right)\geq w>0$
+$\|h\|_{C^{1,\alpha}_{\rm hom,-1}}\leq\varepsilon<\varepsilon_{n},\quad B-D\left(m\left(m+\frac{n}{2}\right)+C_{n}\sup_{\Sigma}\left(f|A|^{2}+f|\nabla A|\right)\right)\geq w>0,$
 $\Gamma\geq\Gamma(n,B,m,w),\text{ and }\gamma_{1}\leq\gamma_{1}(n,\Sigma,F,B,D,m,w),$
@@
 Plugging in the expression, there exists $\tau_{0}$ such that given
-$|h|_{C_{1}}\leq\varepsilon<\varepsilon_{n},\quad B-a\left(C_{n}\sup_{\Sigma}\left(f|A|^{2}+f|\nabla A|\right)\right)-C_{0}\sqrt{a}\geq w>0,\text{ and}$
+$|h|_{C^{\rm hom}_{1;1}}\leq\varepsilon<\varepsilon_{n},\quad B-a\left(C_{n}\sup_{\Sigma}\left(f|A|^{2}+f|\nabla A|\right)\right)-C_{0}\sqrt{a}\geq w>0,\text{ and}$
 $\tau_{0}\geq\tau_{0}(n,\Sigma,F,C_{0},B,\gamma_{2},\varepsilon_{n}),$
@@
 satisfies $(\partial_{t}-Q)u\geq 0$ on its domain if
-$|h|_{C_{1}}\leq\varepsilon<\varepsilon_{n},\quad B-D\left(m(m+\frac{n}{2})+C_{n}\sup_{\Sigma}\left(f|A|^{2}+f|\nabla A|\right)\right)\geq w>0$
+$\|h\|_{C^{1,\alpha}_{\rm hom,-1}}\leq\varepsilon<\varepsilon_{n},\quad B-D\left(m(m+\frac{n}{2})+C_{n}\sup_{\Sigma}\left(f|A|^{2}+f|\nabla A|\right)\right)\geq w>0$
 $\Gamma\geq\Gamma(n,B,m,w),\quad\gamma_{1}\leq\gamma_{1}(n,\Sigma,F,B,D,m,w),$
@@
 Given $\Gamma>0$ and $h$ satisfying (2.9) on $\Sigma\times(0,\tau_{1})$ and $\tau=-\log(1-t)$, assume
-$\|h\|_{C^{1}(\Sigma\times[0,\tau_{1}])}$ $<\varepsilon,$
-$h(0)$ $=\eta\left(\frac{1}{\gamma_{0}e^{\tau_{0}}}f(x)\right)\sum_{i=1}^{m}p_{i}f_{i}\text{ for }|\mathbf{p}|\leq\bar{p}e^{\lambda_{*}\tau_{0}}$
-$\|h(\tau)\|_{L_{W}^{2}(\Sigma)}$ $\leq\mu e^{\lambda_{*}(\tau+\tau_{0})}\text{ for all }\tau\in[0,\tau_{1}],$
-$\text{supp}(\mathcal{Q}_{1}(\cdot,\tau))$$\subseteq\left\{\frac{\Gamma_{0}e^{\frac{(\tau+\tau_{0})}{2}}}{2}\leq f\leq\Gamma_{0}e^{\frac{(\tau+\tau_{0})}{2}}\right\},$
+$\|h\|_{C^{1,\alpha}_{\rm hom,-1}(\Sigma\times[0,\tau_{1}])}$ $<\varepsilon,$
+$h(0)$ $=\eta\left(\frac{1}{\gamma_{0}e^{\tau_{0}}}f(x)\right)\sum_{i=1}^{m}p_{i}f_{i}\text{ for }|\mathbf{p}|\leq\bar{p}e^{-\lambda_{*}\tau_{0}},$
+$\|h(\tau)\|_{L_{W}^{2}(\Sigma)}$ $\leq\mu e^{-\lambda_{*}(\tau+\tau_{0})}\text{ for all }\tau\in[0,\tau_{1}],$
+$\text{supp}(\mathcal{Q}_{1}(\cdot,\tau))$$\subseteq\left\{\frac{\Gamma_{0}e^{(\tau+\tau_{0})}}{4}\leq f\leq\Gamma_{0}e^{(\tau+\tau_{0})}\right\},$
 $|\mathcal{Q}_{1}(\cdot,\tau)|$ $\leq C\left(\frac{1}{\Gamma_{0}e^{(\tau+\tau_{0})}}\right)^{\frac{1}{2}},\text{ and}$
-$|\nabla\mathcal{Q}_{1}(\cdot,\tau)|$ $\leq C\left(\frac{1}{\Gamma_{0}e^{\tau+\tau_{0}}}\right).$
-Then for any $\delta<\varepsilon$, $\Gamma>\Gamma_{0}(n,\Sigma,F)$, $\varepsilon<\varepsilon_{0}(n,\Sigma,F,\Gamma)$, $|\mathbf{p}|<C(n,\Sigma,F,\lambda_{*})$, $\mu<\mu_{0}(n,\Sigma,F,\lambda_{*})$, $\gamma<\gamma_{0}(n,\Sigma,F,\lambda_{*},\delta),$ and $\tau_{0}>C(n,\Sigma,F,\lambda_{*},\Gamma,\bar{p},\varepsilon,\delta)$, there exist $W=W(n,\Sigma,F,\lambda_{*})<\infty$ and $W^{\prime}=W^{\prime}(n,\Sigma,F)<\infty$ such that
-$|h|+|\nabla h|+|\nabla^{2}h|$ $\leq W\left(\frac{|F|^{2}}{4}\right)^{|\lambda_{*}|}e^{\lambda_{*}{(\tau+\tau_{0})}}\text{ for }(x,\tau)\in\Sigma\times[0,\tau_{1}],\text{ and}$
-$|h|+|\nabla h|+|\nabla^{2}h|$ $\leq W^{\prime}\delta\text{ for }(x,\tau)\in\Sigma\times[0,\tau_{1}].$
+$|\nabla\mathcal{Q}_{1}(\cdot,\tau)|$ $\leq C\left(\frac{1}{\Gamma_{0}e^{\tau+\tau_{0}}}\right)$
+for some $\lambda_{*}>0.$ Then for any $\delta<\varepsilon$, $\Gamma>\Gamma_{0}(n,\Sigma,F)$, $\varepsilon<\varepsilon_{0}(n,\Sigma,F,\Gamma)$, $|\mathbf{p}|<C(n,\Sigma,F,\lambda_{*})$, $\mu<\mu_{0}(n,\Sigma,F,\lambda_{*})$, $\gamma<\gamma_{0}(n,\Sigma,F,\lambda_{*},\delta),$ and $\tau_{0}>C(n,\Sigma,F,\lambda_{*},\Gamma,\bar{p},\varepsilon,\delta)$, there exist $W=W(n,\Sigma,F,\lambda_{*})<\infty$ and $W^{\prime}=W^{\prime}(n,\Sigma,F)<\infty$ such that
+$|\nabla^{l}h|$ $\leq W\tilde{r}^{2\lambda_{*}+1-l}e^{-\lambda_{*}{(\tau+\tau_{0})}},\,\,l\in\{0,1,2\}\text{ for }(x,\tau)\in\Sigma\times[0,\tau_{1}],\text{ and}$
+$|\nabla^{l}h|$ $\leq W^{\prime}\delta\tilde{r}^{1-l},\,\,l\in\{0,1,2\}\text{ for }(x,\tau)\in\Sigma\times[0,\tau_{1}].$
 6. Ważewski’s box argument
@@
 Based on Proposition 2.3, we take an $L^{2}_{W}$-orthonormal eigenbasis $\{h_{i}\}_{i\in\mathbb{N}}$ of $L$ on $\Sigma$ and the corresponding eigenvalues $\{\lambda_{i}\}_{i\in\mathbb{N}}$.
-Take $\lambda_{*}<0$ and $m$ such that $\lambda_{m+1}<\lambda_{*}<\lambda_{m}$.
+Take $\lambda_{*}>0$ and $m$ such that $\lambda_{m}<\lambda_{*}<\lambda_{m+1}$.
 We fix a compactly supported bump function $\eta:(0,\infty)\to[0,1]$ such that $\eta(x)=1$ for $x<\frac{1}{2}$ and $\eta(x)=0$ for $x>1$.
 Take any $\mathbf{p}=(p_{1},p_{2},\cdots,p_{m})\in\mathbb{R}^{m},$ and then we can perturb the immersion of $M$ into $\mathbb{R}^{n+1}$ such that
-$F_{\mathbf{p}}(x,0)=F_{M}(x)+\eta\left(\frac{1}{\gamma_{0}e^{\tau_{0}}}f(x)\right)\sum_{i=1}^{m}p_{i}f_{i}\cdot\textbf{n}_{F}(x).$
+$F_{\mathbf{p}}(x,0)=F_{M}(x)+h_{\mathbf{p}}(x)\textbf{n}_{F}(x):=F_{M}(x)+\eta\left(\frac{1}{\gamma_{0}e^{\tau_{0}}}f(x)\right)\sum_{i=1}^{m}p_{i}f_{i}\cdot\textbf{n}_{F}(x).$
 We start the mean curvature flow from $F_{\mathbf{p}}(x,0)$.
@@
 $\hat{F}_{\mathbf{p}}(x,t)=\eta\left(\frac{1}{\Gamma_{0}e^{\tau_{0}}}f(x)\right)F_{\mathbf{p}}(x,t)+\left(1-\eta\left(\frac{1}{\Gamma_{0}e^{\tau_{0}}}f(x)\right)\right)F(x,t),$
-where $F(x,t)$ corresponds to the family of immersions of mean curvature flows of the shrinker $\Sigma$. From the discussion in Section 2.2 we can take $F_{0}=\hat{F}_{\mathbf{p}}$. Then we get
+where $F(x,t)$ corresponds to the family of immersions of mean curvature flows of the shrinker $\Sigma$. From the discussion in Section 2.2, we can take $F_{0}=\hat{F}_{\mathbf{p}}$. Then we get
 $\frac{1}{\sqrt{1-t}}\hat{F}_{\mathbf{p}}(\varphi_{t}(x),t)=F(x)+h_{\mathbf{p}}(x,t)\textbf{n}_{F}(x)$
@@
 In particular, we have
-(6.1) $\left|d\hat{F}_{\mathbf{p}}(\partial_{t}\varphi_{t}(x))\right|\leq\frac{\left|(\partial_{t}\hat{F}_{\mathbf{p}})(\varphi_{t}(x),t)+\frac{1}{2\sqrt{1-t}}F(x)\right|}{\left|\left\langle\textbf{n}_{F}(x),\textbf{n}_{\hat{F}_{\mathbf{p}}}(\varphi_{t}(x))\right\rangle\right|}.$
+$\left|d\hat{F}_{\mathbf{p}}(\partial_{t}\varphi_{t}(x))\right|\leq\frac{\left|(\partial_{t}\hat{F}_{\mathbf{p}})(\varphi_{t}(x),t)+\frac{1}{2\sqrt{1-t}}F(x)\right|}{\left|\left\langle\textbf{n}_{F}(x),\textbf{n}_{\hat{F}_{\mathbf{p}}}(\varphi_{t}(x))\right\rangle\right|}.$
 From (2.9), the function $h=h_{\mathbf{p}}$ satisfies the equation
@@
 In normal coordinates, the second order term in $\mathcal{Q}_{2}$ corresponds to
-(6.2) $\frac{\tilde{h}^{ab}\left(1-|\nabla h|^{2}\right)-\left(|\nabla h|^{2}+\tilde{h}^{cd}\partial_{c}h\,\partial_{d}h\right)g^{ab}}{\left(1+\left\langle\textbf{n}_{F},\widetilde{\textbf{n}}\right\rangle\right)^{2}}\partial_{a}\partial_{b}h\cdot\textbf{n}_{F}.$
+$\frac{\tilde{h}^{ab}\left(1-|\nabla h|^{2}\right)-\left(|\nabla h|^{2}+\tilde{h}^{cd}\partial_{c}h\,\partial_{d}h\right)g^{ab}}{\left(1+\left\langle\textbf{n}_{F},\widetilde{\textbf{n}}\right\rangle\right)^{2}}\partial_{a}\partial_{b}h\cdot\textbf{n}_{F}.$
 Given any $h\in L^{2}_{W}(\Sigma)$, we denote the stable and unstable projections by
@@
 (2) For $\tau\in[0,-\log(1-t_{1})],$
-$\|h_{u}(\cdot,\tau)\|_{L^{2}_{W}}$ $\leq\mu_{u}e^{\lambda_{*}(\tau+\tau_{0})},\text{ and}$
-$\|h_{s}(\cdot,\tau)\|_{L^{2}_{W}}$ $\leq\mu_{s}e^{\lambda_{*}{(\tau+\tau_{0})}}.$
+$\|h_{u}(\cdot,\tau)\|_{L^{2}_{W}}$ $\leq\mu_{u}e^{-\lambda_{*}(\tau+\tau_{0})},\text{ and}$
+$\|h_{s}(\cdot,\tau)\|_{L^{2}_{W}}$ $\leq\mu_{s}e^{-\lambda_{*}{(\tau+\tau_{0})}}.$
 (3) For $\tau\in[0,-\log(1-t_{1})],$
-$|h|_{\bar{g}}$ $\leq\varepsilon_{0},$
+$|h|_{\bar{g}}(x)$ $\leq\varepsilon_{0}\tilde{r},$
 $|\nabla h|_{\bar{g}}$ $\leq\varepsilon_{1},\text{ and}$
-$|\nabla^{2}h|_{\bar{g}}$ $\leq\varepsilon_{2}.$
+$|\nabla^{2}h|_{\bar{g}}$ $\leq\varepsilon_{2}\tilde{r}^{-1}.$
+where $\tilde{r}$ is a fixed smooth function such that $\tilde{r}\geq 1$ and $\tilde{r}=|x|$ for $|x|\geq 2$.
 Our main goal here is to prove that for $\gamma_{0},|\mathbf{p}|<<1$ and $\Gamma_{0},\tau_{0}>>1$, $h_{\mathbf{p}}$ can only leave $\mathcal{B}$ from the unstable side.
 […]
-Given $\gamma_{0},|\mathbf{p}|\leq\bar{p}e^{\lambda_{*}\tau_{0}}<<1$ and $\Gamma_{0},\tau_{0}>>1$, for $h_{\mathbf{p}}$ constructed above, assume $h_{\mathbf{p}}\notin\mathcal{B}[\lambda_{*},\mu_{u},\mu_{s},\varepsilon_{0},\varepsilon_{1},\varepsilon_{2},1].$ If we denote the largest existence time for $F_{\mathbf{p}}(x,\cdot)$ by $T(\mathbf{p})$ and take the largest time $t(\mathbf{p})$ such that $h_{\mathbf{p}}\in\mathcal{B}=\mathcal{B}[\lambda_{*},\mu_{u},\mu_{s},\varepsilon_{0},\varepsilon_{1},\varepsilon_{2},t(\mathbf{p})]$, then we have
-$\|h_{\mathbf{p},u}(\cdot,-\log(1-t(\mathbf{p})))\|_{L^{2}_{W}}=\mu_{u}e^{\lambda_{*}(\tau_{0}-\log(1-t(\mathbf{p})))}.$
+Given $\gamma_{0},|\mathbf{p}|\leq\bar{p}e^{-\lambda_{*}\tau_{0}}<<1$ and $\Gamma_{0},\tau_{0}>>1$, for $h_{\mathbf{p}}$ constructed above, assume $h_{\mathbf{p}}\notin\mathcal{B}[\lambda_{*},\mu_{u},\mu_{s},\varepsilon_{0},\varepsilon_{1},\varepsilon_{2},1].$ If we denote the largest existence time for $F_{\mathbf{p}}(x,\cdot)$ by $T(\mathbf{p})$ and take the largest time $t(\mathbf{p})$ such that $h_{\mathbf{p}}\in\mathcal{B}=\mathcal{B}[\lambda_{*},\mu_{u},\mu_{s},\varepsilon_{0},\varepsilon_{1},\varepsilon_{2},t(\mathbf{p})]$, then we have
+$\|h_{\mathbf{p},u}(\cdot,-\log(1-t(\mathbf{p})))\|_{L^{2}_{W}}=\mu_{u}e^{-\lambda_{*}(\tau_{0}-\log(1-t(\mathbf{p})))}.$
 Proof.
-From the standard ordinary differential equation and parabolic theories, we have that for $|\mathbf{p}|<<1$, we can get $t(\mathbf{p})>0$.
+From the standard ordinary differential equation and parabolic theories, for $|\mathbf{p}|<<1$, we can get $t(\mathbf{p})>0$.
 From the interior estimate and the pseudolocality in [EH, INS], on $M\setminus\left\{f\leq\frac{\Gamma_{0}e^{\tau_{0}}}{2}\right\}$, $F_{\mathbf{p}}(x,t)$ satisfies $|\nabla^{m}A|\leq C_{m}(\frac{1}{\Gamma_{0}e^{\tau_{0}}})^{\frac{m+1}{2}}$ for $t\in[0,2]$.
@@
 $|\nabla(\partial_{t}\hat{F}_{\mathbf{p}}-\Delta^{g_{0}}\hat{F}_{\mathbf{p}})|(x,t)$ $\leq C\left(\frac{1}{\Gamma_{0}e^{\tau_{0}}}\right).$
-Hence, if we denote $\tau=-\log(1-t)$,
-$\text{supp}(\mathcal{Q}_{1}(\cdot,\tau))$ $=\text{supp}\left({\sqrt{1-t}}\left\langle(\partial_{t}\hat{F}_{\mathbf{p}}-\Delta^{g_{0}}\hat{F}_{\mathbf{p}})(\varphi_{t}(\cdot),t),\textbf{n}_{F_{1}}(\cdot)\right\rangle\right)\subseteq\left\{\frac{\Gamma_{0}e^{\frac{(\tau+\tau_{0})}{2}}}{2}\leq f\leq\Gamma_{0}e^{\frac{(\tau+\tau_{0})}{2}}\right\},$
+Hence, if we denote $\tau=-\log(1-t)$, we have
+$\text{supp}(\mathcal{Q}_{1}(\cdot,\tau))$ $=\text{supp}\left({\sqrt{1-t}}\left\langle(\partial_{t}\hat{F}_{\mathbf{p}}-\Delta^{g_{0}}\hat{F}_{\mathbf{p}})(\varphi_{t}(\cdot),t),\textbf{n}_{F_{1}}(\cdot)\right\rangle\right)\subseteq\left\{\frac{\Gamma_{0}e^{(\tau+\tau_{0})}}{2}\leq f\leq\Gamma_{0}e^{(\tau+\tau_{0})}\right\},$
 $|\mathcal{Q}_{1}(\cdot,t)|$ $\leq C\left(\frac{1}{\Gamma_{0}e^{\tau+\tau_{0}}}\right)^{\frac{1}{2}},\text{ and }|\nabla\mathcal{Q}_{1}(\cdot,t)|\leq C\left(\frac{1}{\Gamma_{0}e^{\tau+\tau_{0}}}\right).$
@@
 $\|h\|_{C^{2}(\Sigma\times[0,\tau_{1}])}$ $\leq W^{\prime}\delta<\frac{1}{2}\min\{\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}\},\text{ and}$
-$\|h\|_{C^{2}(\Sigma\times[0,\tau_{1}])}\leq W\left(\frac{|F|^{2}}{4}\right)^{|\lambda_{*}|}e^{\lambda_{*}(\tau+\tau_{0})}.$
+$\|h\|_{C^{2}(\Sigma\times[0,\tau_{1}])}$ $\leq W\left(\frac{|F|^{2}}{4}\right)^{\lambda_{*}}e^{-\lambda_{*}(\tau+\tau_{0})}.$
 From the equation, given $|h|_{C^{2}}<\varepsilon_{0}(n)$,
-$|\mathcal{Q}_{2}|\leq C\left(|h|+|\nabla h|+|\nabla^{2}h|\right)^{2}\leq C\min\left\{W\left(\frac{|F|^{2}}{4}\right)^{|\lambda_{*}|}e^{\lambda_{*}(\tau+\tau_{0})},\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}\right\}.$
+$|\mathcal{Q}_{2}|\leq C\left(|h|+|\nabla h|+|\nabla^{2}h|\right)^{2}\leq C\min\left\{W\left(\frac{|F|^{2}}{4}\right)^{\lambda_{*}}e^{-\lambda_{*}(\tau+\tau_{0})},\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}\right\}.$
 Thus, given $\tau_{0}>>1$,
-(6.3) $\|\mathcal{Q}_{1}\|_{L^{2}_{W}}+\|\mathcal{Q}_{2}\|_{L^{2}_{W}}\leq CW^{2}e^{2\lambda_{*}(\tau+\tau_{0})}\leq Ce^{2\lambda_{*}(\tau+\tau_{0})}.$
-From the definition of $h_{\mathbf{p}}$, the polynomial decay of eigenfunctions (Theorem 2.4), and the exponential decay of the weight, for $|\mathbf{p}|\leq\bar{p}e^{\lambda_{*}\tau_{0}}$, we have
+(6.1) $\|\mathcal{Q}_{1}\|_{L^{2}_{W}}+\|\mathcal{Q}_{2}\|_{L^{2}_{W}}\leq CW^{2}e^{-2\lambda_{*}(\tau+\tau_{0})}\leq Ce^{-2\lambda_{*}(\tau+\tau_{0})}.$
+From the definition of $h_{\mathbf{p}}$, the polynomial decay of eigenfunctions (Theorem 2.4), and the exponential decay of the weight, for $|\mathbf{p}|\leq\bar{p}e^{-\lambda_{*}\tau_{0}}$, we have
 (6.2) $\left|\langle h_{\mathbf{p}}(\cdot,0),h_{j}\rangle_{L^{2}_{W}}-p_{j}\right|$ $\leq Ce^{-\frac{\gamma_{0}}{100}e^{\tau_{0}}}\text{ for }j\leq m,\text{ and}$
@@
 From (6.1) and the equation $\partial_{\tau}h=Lh+\mathcal{Q}_{1}+\mathcal{Q}_{2}$,
-(6.5) $\frac{d}{d\tau}\|e^{-\lambda_{*}(\tau+\tau_{0})}h_{\mathbf{p},u}\|_{L^{2}_{W}}$ $\geq(\lambda_{m}-\lambda_{*})\|e^{-\lambda_{*}(\tau+\tau_{0})}h_{\mathbf{p},u}\|_{L^{2}_{W}}-Ce^{\lambda_{*}\tau_{0}},\text{ and}$
-$\frac{d}{d\tau}\|e^{-\lambda_{*}(\tau+\tau_{0})}h_{\mathbf{p},s}\|_{L^{2}_{W}}$ $\leq(\lambda_{m+1}-\lambda_{*})\|e^{-\lambda_{*}(\tau+\tau_{0})}h_{\mathbf{p},u}\|_{L^{2}_{W}}+Ce^{\lambda_{*}\tau_{0}}.$
+(6.3) $\frac{d}{d\tau}\|e^{\lambda_{*}(\tau+\tau_{0})}h_{\mathbf{p},u}\|_{L^{2}_{W}}$ $\geq(\lambda_{*}-\lambda_{m})\|e^{\lambda_{*}(\tau+\tau_{0})}h_{\mathbf{p},u}\|_{L^{2}_{W}}-Ce^{-\lambda_{*}\tau_{0}},\text{ and}$
+$\frac{d}{d\tau}\|e^{\lambda_{*}(\tau+\tau_{0})}h_{\mathbf{p},s}\|_{L^{2}_{W}}$ $\leq(\lambda_{*}-\lambda_{m+1})\|e^{\lambda_{*}(\tau+\tau_{0})}h_{\mathbf{p},u}\|_{L^{2}_{W}}+Ce^{-\lambda_{*}\tau_{0}}.$
 In particular, given $\tau_{0}>C(n,\mu_{s},\mu_{u}),$
-$\|h_{\mathbf{p},s}(\cdot,-\log(1-t(\mathbf{p})))\|_{L^{2}_{W}}<\mu_{s}e^{\lambda_{*}(\tau_{0}-\log(1-t(\mathbf{p})))},$
+$\|h_{\mathbf{p},s}(\cdot,-\log(1-t(\mathbf{p})))\|_{L^{2}_{W}}<\mu_{s}e^{-\lambda_{*}(\tau_{0}-\log(1-t(\mathbf{p})))},$
 and thus, from the definition of the box, we obtain
-$\|h_{\mathbf{p},u}(\cdot,-\log(1-t(\mathbf{p})))\|_{L^{2}_{W}}=\mu_{u}e^{\lambda_{*}(\tau_{0}-\log(1-t(\mathbf{p})))}.$
+$\|h_{\mathbf{p},u}(\cdot,-\log(1-t(\mathbf{p})))\|_{L^{2}_{W}}=\mu_{u}e^{-\lambda_{*}(\tau_{0}-\log(1-t(\mathbf{p})))}.$
 Moreover, from (6.3), for $\tau>-\log(1-t(\mathbf{p}))$,
-$\|h_{\mathbf{p},u}(\cdot,\tau)\|_{L^{2}_{W}}>\mu_{u}e^{\lambda_{*}(\tau+\tau_{0})}.$
+$\|h_{\mathbf{p},u}(\cdot,\tau)\|_{L^{2}_{W}}>\mu_{u}e^{-\lambda_{*}(\tau+\tau_{0})}.$
 This completes the proof. ∎
@@
 […]
-Given $\gamma_{0},$ $\bar{p}e^{\lambda_{*}\tau_{0}}<<1$ and $\Gamma_{0},\tau_{0}>>1$, there exists $\mathbf{p}\in\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}$ such that
+Given $\gamma_{0},$ $\bar{p}e^{-\lambda_{*}\tau_{0}}<<1$ and $\Gamma_{0},\tau_{0}>>1$, there exists $\mathbf{p}\in\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}$ such that
 (6.4) $h_{\mathbf{p}}\in\mathcal{B}[\lambda_{*},\mu_{u},\mu_{s},\varepsilon_{0},\varepsilon_{1},\varepsilon_{2},1].$
 Proof.
-As in the Ricci flow case, we define $\mathcal{F}:\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}\to\mathbb{R}^{m}$ by
+As in the Ricci flow case, we define $\mathcal{F}:\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}\to\mathbb{R}^{m}$ by
 $\mathcal{F}(\mathbf{p})=\left(\left\langle h_{\mathbf{p}}(t(\mathbf{p})),h_{i}\right\rangle_{L^{2}_{W}}\right)_{i=1,2,\cdots,m}.$
 If we take $\mu_{u}<\frac{\bar{p}}{2}$, from (6.2), the map
-$\mathcal{F}|_{\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}\setminus\overline{B_{2\mu_{u}e^{\lambda_{*}\tau_{0}}}}}\colon\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}\setminus\overline{B_{2\mu_{u}e^{\lambda_{*}\tau_{0}}}}\to\mathbb{R}^{m}\setminus\{0\}$
-is homotopic to the identity through $\mathcal{F}_{s}:=s\mathcal{F}+(1-s){\rm Id}\colon\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}\setminus\overline{B_{2\mu_{u}e^{\lambda_{*}\tau_{0}}}}\to\mathbb{R}^{m}\setminus\{0\}$.
-On the other hand, if
+$\mathcal{F}|_{\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}\setminus B_{2\mu_{u}e^{-\lambda_{*}\tau_{0}}}}\colon\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}\setminus B_{2\mu_{u}e^{-\lambda_{*}\tau_{0}}}\to\mathbb{R}^{m}\setminus\{0\}$
+is homotopic to the identity through $\mathcal{F}_{s}:=s\mathcal{F}+(1-s){\rm Id}\colon\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}\setminus B_{2\mu_{u}e^{-\lambda_{*}\tau_{0}}}\to\mathbb{R}^{m}\setminus\{0\}$.
+We claim that this shows the Proposition. Otherwise, if the proposition is false, then
 $h_{\mathbf{p}}\notin\mathcal{B}[\lambda_{*},\mu_{u},\mu_{s},\varepsilon_{0},\varepsilon_{1},\varepsilon_{2},1]$
-for any $\mathbf{p}\in\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}},$ then Proposition 6 implies that $|\mathcal{F}(\mathbf{p})|>0$ for $\mathbf{p}\in\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}$.
+for all $\mathbf{p}\in\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}$.
+Thus Proposition 6 implies that $|\mathcal{F}(\mathbf{p})|>0$ for all $\mathbf{p}\in\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}$.
+Note that $\mathcal{F}|_{\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}}$ is null-homotopic to a constant map by shrinking $\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}$ to a point and compose it with $\mathcal{F}$.
+Here, we have that the whole homotopy lies in $\mathbb{R}^{m}\setminus\{0\}$ as $|\mathcal{F}(\mathbf{p})|>0$ for all $\mathbf{p}\in\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}$.
 This shows that, in particular,
-$\mathcal{F}|_{\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}\setminus\overline{B_{2\mu_{u}e^{\lambda_{*}\tau_{0}}}}}:\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}\setminus\overline{B_{2\mu_{u}e^{\lambda_{*}\tau_{0}}}}\to\mathbb{R}^{m}\setminus\{0\}$
+$\mathcal{F}|_{\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}\setminus B_{2\mu_{u}e^{-\lambda_{*}\tau_{0}}}}\colon\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}\setminus B_{2\mu_{u}e^{-\lambda_{*}\tau_{0}}}\to\mathbb{R}^{m}\setminus\{0\}$
 is null-homotopic.
-Hence, we derive a contraction. This proves that there exists $\mathbf{p}\in\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}$ such that
+Hence, we derive a contraction. This proves that there exists $\mathbf{p}\in\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}$ such that
 (6.5) $h_{\mathbf{p}}\in\mathcal{B}[\lambda_{*},\mu_{u},\mu_{s},\varepsilon_{0},\varepsilon_{1},\varepsilon_{2},1],$
@@
 […]
-Given an asymptotically conical shrinker $\Sigma^{n}\subseteq\mathbb{R}^{n+1}$, there exists a closed hypersurface $M\subseteq\mathbb{R}^{n+1}$ such that the mean curvature flow $M_{t}$ starting from $M$ develops a type I singularity at time $1$ at the origin modeled on $\Sigma$.
+Given an asymptotically conical shrinker $\Sigma^{n}\subseteq\mathbb{R}^{n+1}$, there exists a closed embedded hypersurface $M\subseteq\mathbb{R}^{n+1}$ such that the mean curvature flow $M_{t}$ starting from $M$ develops a type I singularity at time $1$ at the origin modeled on $\Sigma$.
 Proof.
 For any $\mathbf{p},$ if we blow up the mean curvature flow starting from the corresponding perturbation at $({\bf 0},1)$, the family of rescaled hypersurfaces is the graph of $h_{\mathbf{p}}$ over $\Sigma$.
-From Proposition 6, there exists a vector $\mathbf{p}$ such that $h_{\mathbf{p}}\in\mathcal{B}[\lambda_{*},\mu_{u},\mu_{s},\varepsilon_{0},\varepsilon_{1},\varepsilon_{2},1].$ Based on Theorem 5 and Proposition 4, $h_{\mathbf{p}}$ converges to 00 in $C^{\infty}_{loc}(\Sigma)$.
+From Proposition 6, there exists a vector $\mathbf{p}$ such that $h_{\mathbf{p}}\in\mathcal{B}[\lambda_{*},\mu_{u},\mu_{s},\varepsilon_{0},\varepsilon_{1},\varepsilon_{2},1].$ Based on Theorem 5 and Proposition 4, $h_{\mathbf{p}}$ converges to 00 in $C^{\infty}_{\rm loc}(\Sigma)$.
 The global $C^{2}$ bound of $h_{\mathbf{p}}$ guarantees that the singularity is of type I.
@@
 Take such a shrinker $\Sigma$ and let $\mathcal{C}$ be its asymptotic cone.
-Theorem 1 implies that there exists a smooth closed hypersurface $M\subseteq\mathbb{R}^{n+1}$ such that the level set flow $M_{t}$ starting from $M$ develops a singularity at $({\bf 0},0)\in\mathbb{R}^{n+1}\times\mathbb{R}$ whose tangent flow is modeled on $\Sigma.$ Thus, $M_{0}$ is a hypersurface with a singularity at the origin modeled on the cone $C$ (cf.
+Theorem 1 implies that there exists a smooth closed hypersurface $M\subseteq\mathbb{R}^{n+1}$ such that the level set flow $M_{t}$ starting from $M$ develops a singularity at $({\bf 0},0)\in\mathbb{R}^{n+1}\times\mathbb{R}$ whose tangent flow is modeled on $\Sigma.$ Thus, $M_{0}$ is a hypersurface with a singularity at the origin modeled on the cone $\mathcal{C}$ (cf.
 [CS, Corollary 1.2]).